### v1

# The cost of ensembling: is it always worth combining?

Abstract
Given the continuous increase in dataset sizes and the complexity of forecasting models, the trade-off between forecast accuracy and computational cost is emerging as an extremely relevant topic, especially in the context of ensemble learning for time series forecasting. To asses it, we evaluated ten base models and eight ensemble configurations across two large-scale retail datasets (M5 and VN1), considering both point and probabilistic accuracy under varying retraining frequencies. We showed that ensembles consistently improve forecasting performance, particularly in probabilistic settings. However, these gains come at a substantial computational cost, especially for larger, accuracy-driven ensembles. We found that reducing retraining frequency significantly lowers costs, with minimal impact on accuracy, particularly for point forecasts. Moreover, efficiency-driven ensembles offer a strong balance, achieving competitive accuracy with considerably lower costs compared to accuracy-optimized combinations. Most importantly, small ensembles of two or three models are often sufficient to achieve near-optimal results. These findings provide practical guidelines for deploying scalable and cost-efficient forecasting systems, supporting the broader goals of sustainable AI in forecasting. Overall, this work shows that careful ensemble design and retraining strategy selection can yield accurate, robust, and cost-effective forecasts suitable for real-world applications.
keywords:
Time series, Demand forecasting, Forecasting competitions, Cross-learning, Global models, Forecast combinations , Ensemble learning , Machine learning, Deep learning, Conformal predictions , Green AI
1 Introduction
Ensemble learning has become a powerful and widely adopted approach in time series forecasting due to its ability to improve predictive accuracy and robustness by aggregating the outputs of multiple models. The fundamental idea behind ensemble methods is that combining forecasts from diverse models can help balance their individual strengths and weaknesses, thereby reducing the risk of overfitting and mitigating model-specific biases or errors. This approach leverages the principle that while individual models might make different types of errors, their combination can average out these errors. One of the most intriguing and empirically validated findings in the literature is that simple ensemble strategies, such as taking the mean or median of forecasts, often outperform more sophisticated combination methods like stacking (also known as meta-learning) or weighted ensembles. This phenomenon, known as the forecast combination puzzle (Claeskens et al., 2016), suggests that in many cases the added complexity of estimating optimal weights or training a meta-learner does not yield meaningful gains in accuracy and can even lead to overfitting, especially when the number of forecast points is limited. Indeed, the success of simple combinations is partly due to their ability to remain robust in the presence of model misspecification, parameter uncertainty, and changing data dynamics, issues that more complex ensemble methods may struggle to accommodate effectively.
In general, the benefits of ensemble learning are numerous: it improves forecast accuracy, enhances model stability across time and datasets, and increases resilience to structural breaks or shifts in the data. Ensembles also provide a practical hedge against model uncertainty, allowing practitioners to rely less on the assumption that any single model is correctly specified. In the context of global forecasting models, ensemble methods become especially valuable. Global models are trained across many time series simultaneously, and while they excel at identifying cross-series patterns, they can also suffer from model instability or biases if the chosen architecture fails to generalize across heterogeneous series. Combining multiple global models, each with different biases, architectures, or feature representations, can help counteract these issues by smoothing out individual model errors and reducing the risk of overfitting to shared but spurious signals across series (Wang et al., 2023). Additionally, ensembles serve as a hedge against model selection uncertainty, which is particularly relevant in global settings where the optimal model may vary substantially across groups of series. By pooling predictions, ensembles offer a form of model diversification that improves generalization, making them a powerful and practical tool in modern forecasting systems.
However, ensemble methods are not without drawbacks. One significant limitation is the increased computational cost, especially when combining large numbers of complex models like deep learning architectures or global forecasting models. This cost is magnified when retraining or model tuning is required for each component model at multiple forecast origins. Additionally, from a business perspective, managing and maintaining multiple models adds complexity to the production of forecasts. Moreover, in cases where the component models are very similar or highly correlated, the marginal benefit of ensembling diminishes, potentially offering little improvement over the best individual model (Kolassa, 2011).
Therefore, one of the major advantages of adopting a global modeling approach can be offset when ensemble learning is introduced. Combining several models, even with simple techniques like averaging, inherently increases the computational time, as multiple models must be trained, validated, and maintained, producing substantial computational overhead. In this sense, ensemble learning may weaken one of the core benefits of the global modeling paradigm: reduced training time and resource usage. Moreover, since forecasts are typically generated using cloud computing services with pay-as-you-go pricing models (Fotios Petropoulos & Spiliotis, 2024), increased computational time and resource consumption directly leads to higher forecasting costs for organizations. These costs may also rise exponentially when frequent retraining strategies are adopted. Indeed, it is a widespread practice to update forecasting models whenever new data becomes available, often driven by the belief that frequent updates enhance the model’s ability to adapt to evolving patterns and improve predictive accuracy. However, when ensembles are used, all the models within the ensembles must be retrained to obtain the final predictions, abruptly increasing the costs of forecasting.
This trade-off between performance and computational time in the context of ensemble learning raises important questions about the balance between accuracy gains and cost efficiency, especially in large-scale forecasting applications. However, despite these challenges, ensemble learning remains one of the most widely used strategies in the forecasting industry, which further amplifies concerns about the long-term sustainability and scalability of forecasting systems. Indeed, from a practical perspective, forecast combinations, coupled with frequent retraining, has also significant environmental costs, since the energy consumption of model training extends beyond direct computational costs, contributing to energy consumption and carbon emissions (Schwartz et al., 2020). Therefore, understanding the cost of ensemble forecasting models and the effects of retraining on their performance is of paramount importance for advancing more sustainable forecasting practices.
1.1 Research Question
We aim to address the question ”Are ensembles always the solution?”. Specifically, we focus on the trade-off between accuracy and sustainability, in terms of the cost of producing the forecasts, among different types of ensembles of global forecasting models. This cost is particularly relevant for ensemble models since it becomes exponentially larger with the number of base models used to produce the combined forecasts. To this end, we use ten global models as base learners, ranging from traditional machine learning methods to advanced deep learning architectures. In addition, we evaluate eight ensemble approaches, all trained and tested on the two most recent and comprehensive retail forecasting datasets: the M5 and VN1 competition datasets.
We also investigated the effects of retraining on the forecasting performance of ensembles, compared to that of the base models. In particular, we assess whether reducing the frequency of retraining, by avoiding re-estimation of base models as new data becomes available, can effectively enhance the cost-accuracy trade-off in ensemble forecasting. Hence, we explored a wide range of retraining strategies, from continuous retraining to a no-retraining approach, including various intermediate periodic retraining schemes to encompass the most practical and effective scenarios.
1.2 Contributions
Our contribution is fourfold:
  • 1.
    We provide the first comprehensive study of the cost-accuracy trade-off of ensembles of global models, using 10 distinct methods, some of the most relevant real-world datasets, and evaluating both point and probabilistic predictions.
  • 2.
    We analyze different ensemble strategies to asses the effects of the ensemble size on forecast accuracy and costs.
  • 3.
    We compare various retraining solutions, such as continuous, periodic, and no retraining, across multiple datasets, to quantify the impact of the retraining frequency on the computational cost of ensemble forecasting.
  • 4.
    We provide practical guidelines for organizations and practitioners on assessing when and how frequently to retrain ensemble forecasting models to obtain an effective balance between predictive accuracy and computational cost.
By tackling these issues, this paper contributes to both forecasting and machine learning communities, offering valuable insights into the trade-offs among accuracy, computational efficiency, and sustainability in the use of an ensemble of global forecasting models.
1.3 Overview
The rest of this paper is organized as follows. After a brief review of related works (Section 2), in Section 3 we describe the design of the experiment used in our study. The datasets and their characteristics are presented in 3.1, while the global forecasting models and the ensemble strategies are shown in 3.2. The concepts related to rolling origin evaluation and retrain scenario are explained in 3.3, while the metrics used to assess the accuracy and cost of models are discussed in 3.4. In Section 4 we show the empirical findings of our study, including forecast accuracy, computing time, and cost analysis of the different scenarios and across the different ensembles. Finally, Section 5 contains our summary and conclusions.
2 Related works
The cross-learning approach in time series forecasting has witnessed substantial advancements in recent years, becoming a central theme in contemporary research. Today, global models are commonly used as benchmarks in empirical studies, underscoring their growing relevance in the field Semenoglou et al. (2021). Additionally, theoretical contributions from Montero-Manso & Hyndman (2021) and Montero-Manso (2023) have shown that global models can achieve accuracy comparable to local models with reduced complexity and without relying on assumptions about data similarity. Indeed, global models have proven particularly effective across a variety of forecasting domains, especially in retail forecasting domain (Makridakis et al., 2022a), and several methods have been proposed to further boost their performance (Godahewa et al., 2021a; Bandara et al., 2020, 2021). In particular, ensemble learning is a powerful tool in the forecasters’ hands offering a reliable strategy to enhance accuracy and robustness by combining the strengths of multiple models, not only in the cross-learning context (Wang et al., 2023). Relative to point forecasting, many different combination methods have been proposed: linear combinations with optimal weights (Newbold & Granger, 1974), performance based weighting (Pawlikowski & Chorowska, 2020), criteria-based weighting (Kolassa, 2011), and many different approaches based on stacking, or meta-learning (Wang et al., 2023). However, time series forecasting competitions have shown that simple ensemble strategies, like simple mean or median, are extremely accurate, yet efficient (Makridakis et al., 2022b). Indeed, simple combination schemes are hard to beat, and the simple arithmetic average of predictions with equal weights remains the most widely used and surprisingly effective combination rule (Claeskens et al., 2016).
Most evaluations of global models emphasize point forecast accuracy, likely due to the fact that many machine learning and deep learning algorithms do not natively produce probabilistic outputs (Makridakis et al., 2022c). However, in applications such as supply chain management, quantifying uncertainty is essential, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022). In this regard, the Conformal Inference framework introduced by Vovk et al. (2005) offers a flexible tool for uncertainty quantification, and has been successfully applied to time series forecasting problems (Stankeviciute et al., 2021). Forecast combinations in the context of probabilistic predictions are an active area of research. Different methods have been proposed, from linear pooling, Bayesian model averages, and integral combinations (Wang et al., 2023). However, quantile aggregation via simple average has proven to be effective and accurate (Smyl & Hua, 2019; Busetti, 2017).
From the perspective of model retraining, the most comprehensive investigation in the context of global models has been conducted by Zanotti (2025), who found that reducing the retraining frequency of global models can lower forecasting costs without harming accuracy. Other works, such as Spiliotis & Petropoulos (2024), explored retraining strategies and parameter updates for local exponential smoothing models, while Huber & Stuckenschmidt (2020) examined retraining in retail demand forecasting using a limited range of models and datasets. Although these studies offer useful insights, direct analysis on the effects of retraining ensemble models remains unexplored.
The question of whether ensemble methods are consistently worthwhile is relevant both for the forecasting and the broader machine learning community. Indeed, the growing advocacy for sustainable AI (Schwartz et al., 2020; Getzner et al., 2023) highlights the importance of evaluating the environmental and computational costs of model development and maintenance. Our study directly addresses this gap by systematically evaluating different ensemble strategies’ cost-accuracy trade-off. We also assess how different updating frequencies affect the forecasting accuracy across the ensemble models, with the aim of testing whether less frequent retraining can be an effective tool to manage this trade-off. Answering these questions, this research provides both theoretical clarity and practical guidelines towards more sustainable forecasting practices.
3 Experimental design
This section presents the empirical analysis performed to assess the performance of ensemble models and to determine whether reducing the frequency of retraining can yield forecasting accuracy comparable to that of a baseline scenario, involving continuous retraining, while systematically decreasing the forecasting costs. We begin by introducing the datasets used in the experiments, followed by a description of the ensemble learning models employed. Finally, we outline the evaluation strategy, including the performance metrics, the different retraining scenarios considered, and the approach used to assess forecast performance and costs.
3.1 Datasets
For our experiments, we utilized two prominent retail forecasting datasets: the M5 and VN1 competition datasets. The M5 competition, part of the M-competitions series led by Spyros Makridakis and collaborators, was designed to benchmark forecasting methods in a retail demand setting (Makridakis et al., 2022b). The M5 dataset (Howard & Makridakis, 2020) is widely recognized and extensively studied, comprising 3,049 daily time series representing unit sales of Walmart products. These sales span three main product categories, Food, Hobbies, and Household, across ten stores in three U.S. states: California, Texas, and Wisconsin. Covering the period from 2011 to 2016, the dataset features highly intermittent time series organized hierarchically, enabling multi-level forecasting (e.g., individual SKUs, categories, stores, and states). It also includes relevant exogenous variables such as prices, promotions, and special events like holidays. The VN1 Forecasting – Accuracy Challenge, launched in October 2024 by Flieber, Syrup Tech, and SupChains, marked the first edition of its kind (Vandeput, 2024). This dataset consists of weekly sales data for 15,053 products sold from 2020 to 2024 by various e-vendors, primarily based in the U.S. Unlike the M5 dataset, which contains sales from a single retailer (Walmart) and a limited number of physical stores, the VN1 dataset aggregates sales across 328 warehouses operated by 46 distinct retailers. To the best of our knowledge, we are among the first to benchmark forecasting models on this dataset. Together, these two datasets represent the most recent and comprehensive publicly available collections of time series related to retail demand, providing a strong foundation for generalizing our findings in the domain of demand forecasting.
Table 1: The M5 and the VN1 datasets used in the experiments.
Dataset Frequency N. Series Min Obs per Series
M5 Daily (7) 28.298 730
VN1 Weekly (52) 15.053 157
In both datasets we concentrate on the most granular level (individual SKUs), because the potential gains from reduced retraining are greatest at lower aggregation levels. To ensure a consistent evaluation strategy (Section 3.3), we filtered the series: daily SKUs retained at least two years of data (more than 730 observations), while weekly SKUs required a minimum of three years (more than 157 observations).
3.2 Forecasting models
In this study, we focused exclusively on global forecasting methods, as we used only this category of models as base learners for our ensembles. Global approaches have become standard in many industries dealing with large-scale time series data, such as retail demand forecasting, where predictions must be made for thousands of SKUs (Januschowski et al., 2020). Indeed, our primary objective is to assess the performance of ensemble models obtained by combining the predictions of different global models trained on large datasets.
A global forecasting model can be defined as:
$Y_{i}^{h}=F(\mathcal{Y},\Theta)\mbox{,}$ (1)
where forecasts for a horizon $h$ for each individual time series $Y_{i}$ are generated using a single model $F$ trained on the entire set of time series in a dataset $\mathcal{Y}$. Hence, using the global modeling paradigm, it is possible to leverage cross-learning to allow the model to learn shared patterns across all time series. Indeed, the model parameters $\Theta$ are shared across all series.
Table 2: Global forecasting models used in the experiment.
Machine Learning Deep Learning
Linear Regression (LR) (Godahewa et al., 2021b) MLP (Rosenblatt, 1958)
Random Forest (RF) (Breiman, 2001) LSTM (Hochreiter & Schmidhuber, 1997)
XGBoost (Chen & Guestrin, 2016) TCN (Van den Oord et al., 2016)
LGBM (Ke et al., 2017) NBEATS (Oreshkin et al., 2020)
CatBoost (Prokhorenkova et al., 2018) NHITS (Challu et al., 2022)
To ensure a comprehensive evaluation of the ensemble models, we included both traditional machine learning models and state-of-the-art deep learning techniques as base learners, chosen for their proven effectiveness in time series forecasting and their methodological diversity. In particular, Table 2 shows the ten different global forecasting models we trained in our experiments.
Machine learning models are generally easier to train than deep learning approaches but often rely heavily on extensive and careful feature engineering to achieve strong forecasting performance (Januschowski et al., 2022). In contrast, deep learning models can automatically learn relevant features, such as lags or rolling statistics, within their architecture, reducing the need for manual preprocessing. However, they are typically more challenging to train due to their larger number of hyperparameters, which can significantly influence forecast accuracy (Smyl, 2020). As in Zanotti (2025), we implemented simplified feature engineering pipelines inspired by the top-performing solutions in the M5 and VN1 competitions. Our feature set included standard time series features such as lags, rolling and expanding means, as well as calendar-related variables like year, month, week, and day of the week. We also integrated static metadata, such as store, product, category, and location identifiers, based on the dataset’s characteristics. For the M5 dataset, we further included external variables like special events. Hyperparameters were chosen based on configurations from leading competition entries when available; otherwise, we used the default values recommended by the respective libraries.
3.2.1 Ensemble learning
Ensemble learning aggregates the predictions of multiple base models to enhance forecast accuracy and reliability, especially when different models capture diverse and complementary patterns in the data. This approach is particularly beneficial in global forecasting, where the use of a single model can lead to instability or overfitting due to the heterogeneous nature of the time series being modeled (Wang et al., 2023). To address these challenges, forecast combinations help by balancing out the individual biases and variances of the base models. In a general formulation, an ensemble forecasting model combines the outputs of multiple models. Using the same notation as before, where $\mathcal{Y}$ is the set of all time series and $\Theta^{(j)}$ denotes the parameters of the $j$-th base model $F^{(j)}$, the ensemble prediction for series $Y_{i}$ at horizon $h$ can be defined as:
$Y_{i}^{h}=G\left(F^{(1)}(\mathcal{Y},\Theta^{(1)}),F^{(2)}(\mathcal{Y},\Theta^{(2)}),\dots,F^{(J)}(\mathcal{Y},\Theta^{(J)})\right)\mbox{.}$ (2)
Here, $G$ is the ensemble function (e.g., mean, median, weighted average) that combines the forecasts of $J$ base models.
In our study, we adopted an ensemble approach that combines the models’ predictions using a simple average. The simple average is a widely used method in forecasting to create forecast combinations due to its simplicity, proven effectiveness in improving prediction robustness, and is often more accurate than theoretically optimal combinations (Claeskens et al., 2016). Therefore, being $G$ the simple mean, the ensemble forecast for series $Y_{i}$ at horizon $h$ becomes:
$Y_{i}^{h}=\frac{1}{J}\sum_{j=1}^{J}F^{(j)}(\mathcal{Y},\Theta^{(j)})\mbox{.}$ (3)
Furthermore, since we are not only interested in point forecast accuracy, in the context of probabilistic forecasting, ensemble methods can also be applied to combine predictive distributions. A common and straightforward approach is to average the predicted quantiles across base models (Wang et al., 2023). That is, for each desired quantile level (e.g., 0.1, 0.5, 0.9), the ensemble forecast simply takes the mean of the corresponding quantiles predicted by the individual models. Despite its simplicity, this method is often effective and overall seems to be preferred compared to other combination techniques (Smyl & Hua, 2019). Moreover, equally weighting quantiles through a simple average yields robust and improved forecast skill because the error in estimating optimal weights usually exacerbates the ensemble predictions (Busetti, 2017). Lastly, simple combination methods have the lowest computational time possible 111 Simple combinations, such as the mean or median, require only basic arithmetic operations on the forecast outputs of the base models, avoiding the need for model training or optimization, as is required in more complex approaches like meta-learning. , allowing us to effectively study, without loss of generality, the associated costs of ensembling.
Table 3: Composition of ENSACC and ENSTIME ensembles for the M5 and VN1 datasets.
Dataset Ensemble Models Included
M5 ENSACC (Accuracy-driven combinations)
ENS2A XGBoost, LGBM
ENS3A XGBoost, LGBM, LR
ENS4A XGBoost, LGBM, LR, NBEATSx
ENS5A XGBoost, LGBM, LR, NBEATSx, MLP
ENSTIME (Time-efficient combinations)
ENS2T CatBoost, LR
ENS3T CatBoost, LR, XGBoost
ENS4T CatBoost, LR, XGBoost, MLP
ENS5T CatBoost, LR, XGBoost, MLP, NBEATSx
VN1 ENSACC (Accuracy-driven combinations)
ENS2A MLP, NBEATSx
ENS3A MLP, NBEATSx, NHITS
ENS4A MLP, NBEATSx, NHITS, RF
ENS5A MLP, NBEATSx, NHITS, RF, XGBoost
ENSTIME (Time-efficient combinations)
ENS2T LR, XGBoost
ENS3T LR, XGBoost, CatBoost
ENS4T LR, XGBoost, CatBoost, MLP
ENS5T LR, XGBoost, CatBoost, MLP, TCN
In our study, we employed simple mean ensembles to combine the forecasts of multiple global models, following two distinct selection criteria. The first strategy, which we refer to as ENSACC, combines the base models that demonstrated the highest individual point forecast accuracy (measured by Equation 4). This approach aims to leverage the strengths of top-performing models, under the assumption that their combined output will retain strong predictive performance while potentially offsetting individual weaknesses. The second strategy, ENSTIME, focuses instead on computational efficiency by combining the models with the lowest training and inference times. This ensemble reflects a pragmatic choice for real-world forecasting systems where computational cost is a critical constraint, such as in large-scale retail applications. For both strategies, we constructed ensembles of increasing size, combining the top 2, 3, 4, and 5 models according to each criterion. Table 3 shows the composition of the eight different ensembles for each dataset. This tiered design allows us to assess how forecast accuracy and computational efficiency evolve as additional models are added to the ensemble. Limiting the maximum ensemble size to five models (out of eight or ten, depending on the dataset) reflects a balance between potential accuracy gains and the diminishing returns or increased complexity often observed with larger ensembles (Wang et al., 2023). This systematic approach enables a thorough evaluation of the trade-off between accuracy and computational cost in ensemble-based global forecasting.
All models were implemented in Python using Nixtla’s framework (Nixtla, 2022). Specifically, the mlforecast library was used to train the machine learning models, while neuralforecast was employed for efficiently training the deep learning models.
3.3 Evaluation strategy
In this section, we introduce the concepts of retraining scenarios and rolling window forecasting that we used in our experiment.
3.3.1 Retrain scenario
Following Zanotti (2025), we evaluated multiple retraining scenarios, or retrain windows. A retrain scenario $r$ denotes a positive integer indicating how often the model is updated, or retrained. Specifically, $r$ defines the number of new observations that must be collected before retraining occurs. These scenarios are tailored to the frequency of each dataset, which in turn determines the forecast horizon and business review cycles. Table 4 summarizes the selected retraining windows, test periods, and forecast horizons.
Table 4: The retraining scenarios, the test window, and the horizon for the M5 and the VN1 datasets.
Dataset Frequency Retraining Scenarios (r) Test Window (T) Horizon (h)
M5 Daily (7) 7, 14, 21, 30, 60, 90, 120, 150, 180, 364 364 28
VN1 Weekly (52) 1, 2, 3, 4, 6, 8, 10, 13, 26, 52 52 13
For example, $r=7$ for daily data implies weekly retraining. Each list contains ten scenarios to ensure a broad yet computationally feasible exploration. The scenario with $r=1$, known as continuous retraining, is the most computationally intensive and typically the most accurate, as the model always uses the most recent data. Hence, we adopt it as our benchmark for both accuracy and cost. However, for daily data, we consider $r=7$ (weekly retraining) as the practical benchmark, since daily updates are uncommon in real-world settings. Conversely, the no-retraining scenario $r=T$, where $T$ is the length of the test set, fits the model only once and uses it for the entire forecasting horizon, minimizing computational load but likely yielding the lowest accuracy. All intermediate values $1<r<T$ represent periodic retraining strategies, where both forecast accuracy and computational cost are expected to decrease as $r$ increases.
Since we use global models, the training set at each retraining step includes all time series in the dataset, each extended by $r$ new observations. We evaluate two update strategies: (i) full retraining at each window $r$, or (ii) no update between retrain points, i.e., the model remains unchanged for the next $r$ periods.
3.3.2 Rolling window evaluation
Out-of-sample evaluation is essential in time series forecasting to assess a model’s ability to generalize to unseen data, especially given the potential for structural changes or unanticipated shifts in future values (Tashman, 2000). Among evaluation strategies, rolling origin evaluation is widely recognized as the most appropriate approach (Bergmeir & Benítez, 2012). This method systematically assesses forecast accuracy over multiple iterations by simulating repeated forecasting cycles. The procedure starts by splitting the series into a training and a test set, maintaining the chronological order. At each step, the model is trained on the current training set and used to predict the next $h$ observations. The forecast origin is then shifted forward by a fixed step, and the process is repeated, either with a growing (expanding) or fixed-size training window. Performance metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
Compared to fixed origin evaluation, which offers only a single evaluation point, rolling origin evaluation provides a more robust assessment by capturing performance across varying conditions such as seasonal shifts, level changes, or trend evolutions (Bergmeir & Benítez, 2012). This is particularly beneficial in dynamic settings like retail or supply chain management, where models must adapt to frequent changes. The setup is flexible: the training window can either expand to include all available past data or remain fixed to a specified length. Most practitioners favor the expanding window, especially when time series are short, as it leverages the full data history (Petropoulos & et al., 2022). In our study, we adopt the expanding window strategy to better reflect realistic business forecasting practices and to accommodate short series, such as those in the weekly VN1 dataset. Following Zanotti (2025), Table 4 outlines the key parameters used in our experiments. Lastly, we set the step size to one in all cases to maximize the number of evaluation points across each retraining scenario.
3.4 Evaluation metrics
Evaluating the accuracy of point forecasts in time series analysis remains a debated topic. Although numerous metrics exist to assess model performance, there is no clear consensus in the literature on which metric is superior (Hewamalage et al., 2023). Given that our focus is on SKU-level demand forecasting, where data is typically intermittent, error metrics based on absolute or percentage deviations are suboptimal, as they emphasize the median rather than the full distribution (Kolassa, 2020). Furthermore, due to scale variations across series, employing a scale-independent metric is essential. To address these concerns, following Zanotti (2025), we adopted the Root Mean Squared Scaled Error (RMSSE), introduced by Hyndman & Koehler (2006), as our primary point forecast accuracy measure. RMSSE is calculated as:
$\text{RMSSE}=\sqrt{\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}(y_{t}-\hat{y}_{t})^{2}}{\frac{1}{n-s}\sum_{t=s+1}^{n}(y_{t}-y_{t-s})^{2}}}\mbox{.}$ (4)
This metric compares the mean squared error of a forecast to that of a seasonal naïve benchmark, providing a relative performance measure. It was the official evaluation metric used in the M5 competition (with $s=1$) (Makridakis et al., 2022b), one of the datasets analyzed in our study. Lower RMSSE values indicate more accurate forecasts.
Beyond point accuracy, we also assessed the probabilistic quality of forecasts across different retraining scenarios. Since most of the machine learning and deep learning models we use do not natively produce probabilistic outputs, we employed Conformal Inference to generate prediction intervals. Conformal Inference is a versatile framework that quantifies predictive uncertainty using a calibration (validation) set, without relying on strong distributional assumptions (Vovk et al., 2005). Originally developed for i.i.d. data, it has recently been adapted to time series contexts (Stankeviciute et al., 2021). Its properties, distribution-free, model-agnostic, efficient, and suitable for small datasets, make it especially well-suited for forecasting benchmarks involving diverse models and datasets. As in Zanotti (2025), to evaluate the quality of the prediction intervals, we used the Quantile Loss (QL), also known as Pinball Loss, and its average form, the Multi-Quantile Loss (MQL):
$\text{QL}=\frac{1}{h}\sum_{t=n+1}^{n+h}\left(q\cdot(y_{t}-\hat{y}_{t})\cdot\mathbb{I}_{y_{t}\geq\hat{y}_{t}}+(1-q)\cdot(\hat{y}_{t}-y_{t})\cdot\mathbb{I}_{y_{t}<\hat{y}_{t}}\right)\mbox{,}$ (5)
$\text{MQL}=\frac{1}{\mathcal{Q}}\sum_{q\in\mathcal{Q}}\text{QL}(q)\mbox{.}$ (6)
As a proper scoring rule, QL enables a rigorous evaluation of probabilistic forecasts (Kolassa, 2016). MQL, particularly in its weighted form, was the principal metric in the M5 Uncertainty competition (Makridakis et al., 2022c).
For our analysis, we constructed prediction intervals around the median and six central levels: 60%, 70%, 80%, 90%, 95%, and 99%, yielding a total of 13 quantiles. The lower quantiles (e.g., median, 60%, 70%) describe the forecast center, while the upper quantiles (90% and above) are crucial for assessing tail risks, key to managing safety stock in retail settings (Barrow & Kourentzes, 2016). These quantiles offer a well-rounded view of forecast uncertainty. To ensure valid estimation of predictive uncertainty, conformal prediction intervals were computed on validation sets at least twice the length of the forecast horizon. This requirement limited the number of time series we could retain in the dataset, as explained in previous sections.
In addition to accuracy, we explicitly evaluated the computational cost of generating forecasts by measuring Computing Time (CT), defined as the number of seconds needed to train a model and produce $h$ step-ahead forecasts (Zanotti, 2025). CT directly reflects forecasting costs in real-world deployments, especially when using cloud-based infrastructure with pay-as-you-go pricing (Spiliotis & Petropoulos, 2024). Thus, CT was recorded for each model across all retraining scenarios, with lower CT values indicating more efficient forecasting.
To facilitate consistent comparisons, results for all evaluation metrics were normalized relative to the baseline (most frequent) retraining scenario for each dataset frequency. To statistically assess performance differences across scenarios, we applied the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).
All experiments were conducted using a Microsoft Azure NC6s_v3 cloud instance running Ubuntu 24, equipped with 1 GPU, 6 CPU cores, and 112 GB of memory.
4 Results and discussion
This section presents and interprets the empirical findings of our study, highlighting the interplay between accuracy, probabilistic performance, computing time, and total cost across different retraining strategies and ensemble configurations. We draw insights from both the M5 and VN1 datasets to identify consistent patterns and dataset-specific nuances.
Table 5: Overall forecasting performance and computational cost for each method across datasets. RMSSE, MQL, and CT (in seconds). Minimum values per column are highlighted in bold.
Type Method M5 VN1
RMSSE MQL CT RMSSE MQL CT
ML LR 0.777 0.267 11373 6.549 2.896 236
RF 1.868 2.590 24862
XGBoost 0.755 0.258 15417 1.890 2.469 530
LGBM 0.771 0.256 44429 3.542 2.625 7824
CatBoost 0.947 0.263 10424 5.762 2.845 805
DL MLP 0.821 0.281 17584 1.543 2.492 962
LSTM 1.913 2.843 1284
TCN 0.865 0.290 33364 1.913 2.843 1127
NBEATSx 0.815 0.279 21226 1.698 2.626 1244
NHITS 0.828 0.284 21969 1.699 2.632 1251
ENSACC Ens2A 0.757 0.255 59846 1.472 2.369 2205
Ens3A 0.757 0.256 71219 1.517 2.410 3456
Ens4A 0.758 0.249 92445 1.524 2.386 28318
Ens5A 0.763 0.251 110029 1.544 2.375 28849
ENSTIME Ens2T 0.814 0.230 21797 5.719 2.545 766
Ens3T 0.776 0.223 37214 5.536 2.588 1572
Ens4T 0.764 0.218 54797 4.472 2.419 2533
Ens5T 0.763 0.220 76024 3.672 2.339 3660
Table 5 summarizes the performance and computational time of all forecasting methods evaluated in this study, across both the M5 and VN1 datasets. The models are grouped into four categories: Machine Learning (ML), Deep Learning (DL), ENSACC (ensembles optimized for accuracy), and ENSTIME (ensembles optimized for computational efficiency). For each method, the table reports three key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to assess point forecast accuracy, MQL (Multi-Quantile Loss) to evaluate probabilistic forecast performance, and CT (Computing Time, in seconds) to reflect the computational time under a cloud-based setting. Overall 222 The overall results are derived from the baseline scenario, that is $r=7$ for M5 and $r=1$ for VN1, as this setting is regarded as standard in both theoretical and practical applications. , the models evaluated achieved better absolute performance on the M5 dataset than on VN1. Several factors may explain this difference, including the larger dataset size, the higher frequency of the time series in M5, the availability of rich external regressors (e.g., promotions, special events), and the presence of well-established benchmark hyperparameter settings, many of which were not available for VN1 during model training.
Analyzing the point forecasting accuracy, the RMSSE confirms the effectiveness of ensemble methods in both datasets. In general, ENSACC ensembles (those optimized for accuracy) consistently outperform individual base models. The accuracy improvements, however, show diminishing returns beyond three or four models, aligning with the well-documented phenomenon in ensemble literature where the marginal gain from adding more models decreases. Notably, while ENSACC ensembles dominate in accuracy, ENSTIME ensembles (optimized for computational efficiency) still provide competitive results in the M5 dataset, demonstrating that even time-efficient combinations can retain good predictive power. Interestingly, ENSTIME ensembles show increasing returns in accuracy as more models are added: their performance consistently improves with each additional component. Indeed, in contrast to ENSACC ensembles, where combining just two models is often sufficient to achieve optimal performance (as illustrated in Figure 1), ENSTIME combinations benefit from larger ensemble sizes, with forecast accuracy continuing to improve as more low-cost models are included.
Figure 1: Friedman-Nemenyi test results of ensembles based on RMSSE.
A similar pattern emerges for probabilistic accuracy evaluated using Multi-Quantile Loss (MQL). Compared to base models, ensembles achieve superior uncertainty quantification, benefiting from the diversity in quantile predictions among different models. Simple averaging across quantiles proves to be a highly effective strategy. Also, in this case, ENSACC forecast combinations exhibit diminishing returns in accuracy as more models are added. On the contrary, increasing the number of models in the ensemble based on computational efficiency leads to consistently improved probabilistic forecasting performance, which may even outperform the ENSACC combinations. As shown in Figure 2, ENSTIME ensembles achieve significantly better probabilistic forecasting performance on the M5 dataset compared to all accuracy-based combinations. In contrast, for ENSACC ensembles, combining more than two base models does not yield additional improvements, confirming once again that a small, high-performing subset is sufficient in that case.
Figure 2: Friedman-Nemenyi test results of ensembles based on MQL.
Unsurprisingly, ensemble models require significantly more computing time than individual models, with algorithms like LGBM and RF contributing disproportionately to this overhead. ENSACC ensembles are especially time-intensive, as they often include the most computationally demanding models. In contrast, ENSTIME ensembles maintain a more manageable runtime profile by design, though they still incur a higher computational cost than their individual components due to the cumulative effect of combining multiple models.
Figure 3 presents the point forecast accuracy of each ensemble model across various retraining scenarios for both the M5 and VN1 datasets. To facilitate comparison, the results are expressed in relative terms with respect to the baseline retraining scenarios ($r=7$ for M5 and $r=1$ for VN1). The RMSSE trajectories of the ENSACC ensembles exhibit remarkable stability across retraining periods. In the M5 dataset, their accuracy remains virtually unchanged regardless of retraining frequency, while in the VN1 dataset, performance even improves in some scenarios. For most ensemble models, particularly under low-frequency retraining (i.e., periodic updates), performance is nearly indistinguishable from the baseline. Although slight accuracy degradation is observed for the ENSTIME ensembles under less frequent retraining, it remains modest—below 2% in M5 and under 5% in VN1, even in the no-retraining condition. These findings build upon and extend the results of Zanotti (2025), suggesting that reducing retraining frequency does not significantly harm the point forecast accuracy of global model ensembles. This robustness likely stems from the relative stability of the datasets, which do not exhibit substantial structural changes or concept drift. Under such conditions, base models continue to capture underlying demand patterns effectively over time, and ensemble combinations further enhance stability as retraining becomes less frequent.
Figure 3: RMSSE results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Similarly, Figure 4 illustrates the relative accuracy of the ensemble models in a probabilistic forecasting context, as measured by the Multi-Quantile Loss (MQL). For the M5 dataset, we observe a clear trend where probabilistic accuracy decreases as retraining becomes less frequent, indicating that regular updates are beneficial for maintaining high-quality uncertainty estimates. This pattern holds across the different ensemble strategies. However, for ENSACC ensembles, the decline in accuracy is minimal for frequent retraining scenarios and becomes slightly more noticeable at higher levels of retraining, though it remains within 5 percentage points. A consistent performance gap is also evident between ENSACC and ENSTIME ensembles, with the former outperforming the latter as the retraining scenario increases. This difference is largely attributable to the inclusion of CatBoost in the ENSTIME ensembles, a model that shows relatively weak performance in this setting. In contrast, the VN1 dataset reveals a different behavior. Here, the relationship between retraining frequency and probabilistic accuracy follows a near-convex pattern. Performance initially improves with less frequent retraining, peaking around $r=4$, and begins to decline thereafter. This suggests that very frequent updates may not be necessary for probabilistic performance in stable, lower-frequency datasets like VN1. Nevertheless, ENSACC ensembles continue to outperform ENSTIME ones across the retraining scenarios. Interestingly, the rate of performance deterioration with increased retraining window is comparable across both ensemble strategies in VN1, indicating that the marginal impact of retraining frequency is relatively uniform across the proposed ensemble techniques. Also these results on the probabilistic performance of ensembles are comparable to those obtained by Zanotti (2025) on the base models.
Figure 4: MQL results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
The Friedman-Nemenyi tests, used to compare the accuracy produced by the same model over different retraining scenarios, confirm the previous results (see Supplementary material). Specifically, in the evaluation of both point forecast accuracy and probabilistic forecast accuracy, some level of periodic retraining is at least as good as the continuous retraining scenario.
Figure 5 illustrates the relative computational time (CT) across different retraining scenarios for both the M5 and VN1 datasets. As anticipated, CT decreases sharply as the retraining interval increases, with the reduction following an approximately exponential pattern. This effect is more pronounced in the VN1 dataset, likely due to its smaller size compared to M5. Specifically, moving from the baseline retraining scenario to the first periodic setting ($r=14$) results in a roughly 33% reduction in computing time for M5, while for VN1, CT is nearly halved.
Figure 5: CT results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
As dataset size increases, the marginal gains in CT from reducing retraining frequency appear to depend more on the number of models in the ensemble (ensemble depth) than on the ensemble type (ENSACC vs. ENSTIME). Simpler ensembles composed of only 2 or 3 models benefit most from reduced retraining, exhibiting steeper declines in CT. Moreover, ENSACC and ENSTIME ensembles show very similar CT profiles when matched by depth, suggesting that computational efficiency in this context is more a function of ensemble size than selection strategy. This distinction is less pronounced in the VN1 dataset, given its smaller scale and shorter time series. These findings have direct implications for forecasting cost management, emphasizing the importance of considering retraining frequency and ensemble size when selecting models for large-scale operational use.
Overall, the evidence from Figures 3, 4, and 5, along with the statistical significance tests, indicates that reducing the retraining frequency of ensemble models does not negatively impact their predictive accuracy. At the same time, it lowers the computational time required to generate forecasts. However, effectively leveraging this time reduction depends on carefully considering the ensemble size. As more models are included in the ensemble, the relative benefit of periodic retraining diminishes compared to using only a single model as in Zanotti (2025). Consequently, both the retraining frequency and the ensemble depth play a critical role in managing computational efficiency, and ultimately the cost, of ensemble-based forecasting systems.
As previously noted, computational time can be directly translated into monetary cost for organizations. In line with Nikolopoulos & Petropoulos (2018); Fotios Petropoulos & Spiliotis (2024); Zanotti (2025), we adopted standard pricing assumptions for cloud computing services to estimate both the overall cost of forecasting and the specific costs associated with each retraining scenario for the base global models and each ensemble 333 It is important to note that the costs have been normalized by the number of SKUs in each dataset, allowing for direct comparison across datasets. This normalization enables meaningful conclusions to be drawn regarding the impact of time series frequency on forecasting costs. . Figures 6, 7, 8 9, and 10 illustrate the estimated forecasting costs for a large-scale retailer, assuming a computing service rate of \$3.5 per hour, with 200,000 unique SKUs and 5,000 stores. Total forecast production cost, computed by combining training time and forecast generation under a cloud-computing pricing model, highlights the trade-off central to our study. ENSACC ensembles incur the highest costs, rapidly reaching millions of dollars under frequent retraining. In contrast, ENSTIME ensembles offer a more sustainable alternative, reducing costs by over 50% in many cases. Moreover, as dataset size increases, the trade-off between forecast accuracy and cost becomes significantly more pronounced. This is clearly illustrated in Figures 7 and 8. For the VN1 dataset, ensemble methods offer noticeable improvements in both point and probabilistic forecasting accuracy compared to most individual base models, while incurring only a modest increase in cost. In this context, there is little reason to favor ENSTIME ensembles over ENSACC ones, as the cost difference is minor and the performance gain from accuracy-based combinations is evident. However, the situation is quite different for the M5 dataset. Here, ensemble models, particularly those optimized for accuracy, are significantly more expensive than individual models. Additionally, the performance advantage of ENSACC ensembles over the most accurate base models is minimal in terms of both RMSSE and MQL. In contrast, ENSTIME ensembles can achieve comparable levels of accuracy with substantially lower additional cost, making them a more attractive option. It is also clear that increasing the number of models in an ensemble does not necessarily lead to proportional gains in forecasting accuracy, which raises questions about the cost-effectiveness of larger combinations. Notably, forecast combinations that disregard the computational efficiency of their base models can lead to excessive and unjustified costs. This is particularly evident in the case of Ens4A and Ens5A on the VN1 dataset, where the inclusion of Random Forest (a highly expensive model in this context) results in disproportionately high forecasting costs with only marginal accuracy improvements.
Figure 6: Cost analysis. Overall cost results for the M5 and VN1 datasets.
Figure 7: Cost analysis. RMSSE vs Cost (\$) for the M5 and VN1 datasets.
Figure 8: Cost analysis. MQL vs Cost (\$) for the M5 and VN1 datasets.
Finally, 9 and 10 show the comparison between base models and ensembles across all retraining scenarios. As expected, forecasting costs decrease exponentially with less frequent retraining. Across both datasets, ENSTIME ensembles consistently incur lower costs than their ENSACC counterparts. Moreover, the cost-saving benefits of reduced retraining frequency become more pronounced as dataset size increases. In the VN1 dataset, cost differences between retraining scenarios remain relatively small, even for ensemble models, suggesting limited sensitivity to retraining frequency. In contrast, the M5 dataset exhibits steady, incremental reductions in cost as retraining becomes less frequent. For example, in the M5 setting, the average cost of continuous retraining for ensemble models is approximately \$2,250,000, more than triple the average cost of individual base models. When retraining is entirely eliminated, this figure drops to around \$500,000, representing a cost reduction of over 75%. However, even in the no-retraining scenario, ensemble models remain roughly twice as expensive as the average base model. Interestingly, the cost gap between ENSACC and ENSTIME ensembles narrows for smaller ensembles (e.g., those with two or three base models) as retraining frequency decreases. This indicates that with less frequent retraining, accuracy-based combinations (ENSACC) become increasingly comparable to efficiency-based ones (ENSTIME) in terms of forecasting cost. Thus, infrequent retraining can help align performance-optimized ensemble strategies with more budget-conscious forecasting objectives.
Figure 9: Cost analysis. Comparison between base models and ensembles for the M5 dataset. Each method and retrain scenario combination are shown in relative terms with respect to the baseline scenario, $r=7$.
Figure 10: Cost analysis. Comparison between base models and ensembles for the VN1 dataset. Each method and retrain scenario combination are shown in relative terms with respect to the baseline scenario, $r=1$.
One might argue that, for a large retailer, these costs (and the potential savings) are relatively minor. However, it’s important to emphasize that the cost reductions achieved through a combination of less frequent retraining, smaller ensemble sizes, and more efficient model combinations typically come with no loss in forecasting accuracy, particularly for point forecasts. This makes such strategies both economically and operationally attractive. These results further support our findings. ENSACC ensembles dominate in accuracy but at a significant computational premium. ENSTIME ensembles represent a middle ground. Importantly, ensembles of just two or three models (e.g., ENS2A, ENS3T) often achieve near-optimal performance, providing a compelling trade-off between complexity, cost, and accuracy, and lowering the retraining frequency can be a good strategy to improve this trade-off even in the context of ensemble learning. This comprehensive analysis reinforces the core message of the article: ensemble methods offer strong accuracy benefits, but their cost can be substantial, especially under frequent retraining. Balancing accuracy and sustainability requires a deliberate model selection and retraining design, with lightweight ensembles, like ENSTIME, being a possible solution in large-scale forecasting systems.
5 Conclusions
This study investigated the cost-effectiveness of ensemble learning in global time series forecasting, focusing on the interplay between forecast accuracy, computational efficiency, and retraining strategies. We evaluated ten global forecasting models, spanning both classical machine learning and deep learning techniques, alongside eight ensemble configurations, across two large-scale and industry-relevant datasets: the M5 and VN1 retail forecasting datasets. By systematically comparing different ensemble sizes, selection criteria (accuracy-based vs. efficiency-based), and retraining frequencies, our aim was to assess whether the accuracy gains of ensemble methods justify their additional computational cost, particularly in large-scale operational settings.
Our findings showed that ensemble models consistently improve both point and probabilistic forecasting performance over most individual base models. This improvement is also relevant for probabilistic accuracy, where ensembles provide better uncertainty quantification through the aggregation of diverse quantile predictions. The results support the notion that ensemble learning offers a robust mechanism for improving predictive reliability and mitigating individual model weaknesses, even in the context of global modeling. However, these benefits come at a cost. Ensemble methods are substantially more computationally intensive than single-model approaches, particularly when composed of complex or resource-heavy algorithms such as LightGBM or Random Forest. These computational demands translate directly into higher forecasting costs under cloud-based computing environments.
Our analysis further distinguished between two ensemble design philosophies: ENSACC (accuracy-based) and ENSTIME (efficiency-based). While ENSACC ensembles generally deliver higher accuracy, especially when composed of two or three strong base models, they are also more expensive. In contrast, ENSTIME ensembles, composed of lightweight, computationally efficient models, can achieve comparable accuracy at a much lower cost, making them an attractive compromise in budget-sensitive applications. Most importantly, we find that increasing the number of models in an ensemble does not always yield proportional accuracy gains and often results in rapidly escalating costs. Small ensembles (with two or three base models) are often sufficient to capture most of the benefit in accuracy.
Crucially, our study explored the effects of retraining frequency on ensemble performance and cost. Even if these effects are more pronounced for the base models alone, we showed that less frequent retraining can reduce computational time, without compromising point forecast accuracy, also in the context of ensembles. In many scenarios, especially for relatively stable datasets, even the no-retraining setup performed nearly as well as continuous retraining, confirming and extending the findings from previous work on global models (Zanotti, 2025). Probabilistic forecasting accuracy is more sensitive to retraining, especially for higher-frequency datasets like M5, but performance declines remain modest for moderate retraining intervals. These insights are highly relevant for real-world forecasting systems, where minimizing operational cost without sacrificing accuracy is essential. Indeed, organizations can achieve economic benefits by optimizing their retraining strategies without compromising the forecast quality.
From a practical standpoint, our findings offer several actionable guidelines for forecasters and organizations deploying large-scale forecasting systems. First, ensembles should be kept small and strategically designed, prioritizing accuracy or efficiency depending on the business context. Second, frequent retraining is often unnecessary. Overall, monthly retraining emerges as a practical compromise for balancing probabilistic forecast accuracy with computational costs. However, if the primary forecasting objective is point prediction, even less frequent retraining intervals can be adopted without significantly affecting performance. Third, ENSTIME strategies represent a viable, low-cost pathway to ensemble forecasting, especially when forecast robustness is desired without substantial computational investment. Retailers and other large-scale forecasters can leverage these insights to build forecasting systems that are both accurate and sustainable. Indeed, these findings also carry broader implications for the sustainability of cross-learning driven forecasting systems. Reducing the frequency of retraining not only lowers operational costs but also leads to significant energy savings, thereby enhancing the environmental sustainability of forecasting processes. This aligns with the principles of ”Green AI,” which advocates for the responsible use of computational resources to minimize the ecological footprint of machine learning applications.
While our findings provide strong evidence on the effects of ensembling global forecasting models, some limitations remain. First, we restricted our experiments to two retail datasets, which, while comprehensive and realistic, may not capture the full diversity of time series behaviors found in other domains such as finance, health care, or energy. Additionally, this study operates under the assumption that the data-generating process remains stable, without notable trends or concept drift. However, in many real-world scenarios, this assumption may not hold, potentially affecting the reliability of less frequent retraining strategies. Moreover, our ensemble strategies used simple averaging without dynamic weighting or more sophisticated stacking techniques, which may further improve performance at the cost of added complexity and computation time. We also assumed static cost estimates for computational resources, whereas real-world costs may vary based on cloud infrastructure, parallelism, or vendor-specific pricing. We also deliberately focused on a global modeling approach only, but local or hybrid frameworks may offer different cost-accuracy trade-offs worth exploring. Finally, we encourage further studies to explore the possible trade-off between cost and forecast stability in the context of ensemble learning. In fact, the stability of forecasts is a highly relevant yet underexplored topic, with important implications for model selection in time series forecasting.
In conclusion, our study reinforces the value of ensemble forecasting in global modeling, showing that well-designed ensemble strategies, particularly those leveraging periodic retraining and lightweight components, can deliver high accuracy at a fraction of the computational cost. Striking the right balance between accuracy, efficiency, and retraining frequency is key to deploying scalable and sustainable forecasting systems in practice.
Supplementary material
In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets.
The Tables 6 and 7 show the forecast accuracy of the different ensembles along the examined retrain scenarios for the M5 daily dataset, while Table 8 depicts the computing time in seconds.
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757
Ens3A 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757
Ens4A 0.758 0.758 0.758 0.759 0.758 0.758 0.758 0.758 0.759 0.759
Ens5A 0.763 0.763 0.763 0.764 0.763 0.764 0.763 0.764 0.764 0.764
Ens2T 0.814 0.827 0.814 0.826 0.822 0.830 0.826 0.827 0.828 0.824
Ens3T 0.776 0.786 0.776 0.786 0.784 0.787 0.785 0.786 0.787 0.785
Ens4T 0.764 0.773 0.764 0.773 0.772 0.774 0.772 0.773 0.773 0.772
Ens5T 0.763 0.771 0.762 0.772 0.771 0.772 0.770 0.771 0.772 0.771
Table 6: M5 RMSSE values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 0.255 0.255 0.256 0.256 0.258 0.258 0.259 0.259 0.259 0.262
Ens3A 0.256 0.256 0.257 0.257 0.259 0.259 0.260 0.260 0.261 0.263
Ens4A 0.249 0.250 0.251 0.251 0.253 0.254 0.255 0.255 0.256 0.259
Ens5A 0.251 0.251 0.252 0.253 0.255 0.256 0.257 0.257 0.259 0.261
Ens2T 0.230 0.265 0.228 0.267 0.267 0.269 0.268 0.269 0.269 0.270
Ens3T 0.223 0.257 0.221 0.258 0.260 0.261 0.261 0.261 0.262 0.263
Ens4T 0.218 0.253 0.216 0.255 0.256 0.257 0.257 0.258 0.259 0.261
Ens5T 0.220 0.254 0.218 0.257 0.258 0.259 0.260 0.260 0.262 0.264
Table 7: M5 MQL values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 59846 32015 23124 17659 10899 8679 7551 7717 6374 5417
Ens3A 71219 38084 27513 20935 12911 10197 8962 9076 7507 6382
Ens4A 92445 54416 42478 34519 25299 22166 20632 20758 18966 17648
Ens5A 110029 68460 55343 46738 36710 33322 31713 31730 29678 28287
Ens2T 21797 11591 8297 6178 3705 2803 2494 2472 2025 1674
Ens3T 37214 19780 14141 10591 6272 4826 4243 4223 3427 2834
Ens4T 54797 33824 27005 22811 17684 15982 15323 15196 14139 13473
Ens5T 76024 50156 41971 36394 30071 27951 26994 26878 25598 24739
Table 8: M5 CT values for each method and retrain scenario combination.
The Tables 9 and 10 show the forecast accuracy of the different ensembles along the examined retrain scenarios for the VN1 weekly dataset, while Table 11 depicts the computing time in seconds.
Method 1 2 3 4 6 8 10 13 26 52
Ens2A 1.472 1.479 1.480 1.453 1.444 1.443 1.447 1.431 1.428 1.430
Ens3A 1.517 1.520 1.504 1.462 1.451 1.435 1.450 1.429 1.429 1.430
Ens4A 1.524 1.528 1.515 1.472 1.461 1.445 1.459 1.436 1.435 1.436
Ens5A 1.544 1.543 1.541 1.495 1.484 1.468 1.480 1.459 1.459 1.455
Ens2T 5.719 5.687 5.778 5.701 5.671 5.696 5.639 5.708 5.797 5.939
Ens3T 5.536 5.492 5.556 5.400 5.326 5.369 5.112 5.246 5.497 5.782
Ens4T 4.472 4.426 4.643 4.296 4.256 4.310 4.162 4.159 4.412 4.659
Ens5T 3.672 3.615 3.807 3.513 3.481 3.523 3.410 3.406 3.603 3.798
Table 9: VN1 RMSSE values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
Ens2A 2.369 2.371 2.370 2.326 2.343 2.358 2.356 2.360 2.405 2.467
Ens3A 2.410 2.399 2.399 2.338 2.335 2.353 2.351 2.362 2.411 2.459
Ens4A 2.386 2.399 2.376 2.326 2.341 2.355 2.349 2.344 2.402 2.440
Ens5A 2.375 2.371 2.383 2.318 2.334 2.341 2.357 2.346 2.405 2.446
Ens2T 2.545 2.551 2.575 2.576 2.594 2.601 2.601 2.608 2.660 2.694
Ens3T 2.588 2.582 2.652 2.595 2.613 2.634 2.617 2.591 2.660 2.725
Ens4T 2.419 2.431 2.473 2.439 2.440 2.466 2.463 2.446 2.532 2.572
Ens5T 2.339 2.335 2.355 2.341 2.361 2.380 2.391 2.380 2.456 2.492
Table 10: VN1 MQL values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
Ens2A 2205 1228 940 748 607 508 460 460 361 312
Ens3A 3456 1932 1477 1175 952 796 722 721 569 492
Ens4A 28318 14312 10192 7352 5307 3880 3183 3291 1831 1094
Ens5A 28849 14586 10391 7498 5414 3962 3252 3361 1876 1125
Ens2T 766 395 287 212 157 120 102 104 67 48
Ens3T 1572 807 574 419 306 229 192 197 118 79
Ens4T 2533 1341 983 747 571 452 394 398 277 217
Ens5T 3660 1980 1477 1141 893 724 641 652 477 392
Table 11: VN1 CT values for each method and retrain scenario combination.
Figures 11 and 12 show the results of the Friedman-Nemenyi test in the context of point forecast accuracy for the M5 and VN1 datasets respectively. Figures 13 and 14 present the test results related to the probabilistic evaluation.
Figure 11: M5 Friedman-Nemenyi test results based on RMSSE.
Figure 12: VN1 Friedman-Nemenyi test results based on RMSSE.
Figure 13: M5 Friedman-Nemenyi test results based on MQL.
Figure 14: VN1 Friedman-Nemenyi test results based on MQL.
The costs tables 12 and 13 show the estimated cost in real values of each scenario for M5 daily data.
Method 7 14 21 30 60 90 120 150 180 364 LR 390,732 208,499 150,791 112,547 69,132 52,131 48,474 46,678 38,937 33,151 XGBoost 529,679 281,358 200,778 151,635 88,200 69,495 60,082 60,155 48,189 39,861 LGBM 1,526,424 818,569 593,676 455,075 286,262 228,698 199,337 204,972 170,802 146,252 CatBoost 358,123 189,714 134,258 99,699 58,167 44,164 37,206 38,266 30,622 24,362 MLP 604,120 482,505 441,981 419,823 392,046 383,287 380,697 376,968 368,023 365,525 TCN 1,146,256 960,626 890,924 862,432 815,006 801,612 791,801 791,348 784,201 776,181 NBEATSx 729,263 561,105 514,166 466,683 425,593 411,232 400,953 401,352 393,673 387,058 NHITS 754,783 573,233 510,655 470,077 425,684 409,844 402,731 402,394 393,071 385,470 Average 754,922 509,451 429,654 379,746 320,011 300,058 290,160 290,267 278,440 269,733
Table 12: M5 estimated costs for each base method and retrain scenario combination (in \$).
Method 7 14 21 30 60 90 120 150 180 364 Ens2A 2,056,103 1,099,927 794,454 606,710 374,461 298,193 259,419 265,127 218,991 186,113 Ens3A 2,446,835 1,308,426 945,244 719,256 443,594 350,323 307,893 311,805 257,928 219,264 Ens4A 3,176,097 1,869,531 1,459,411 1,185,940 869,187 761,555 708,845 713,157 651,601 606,322 Ens5A 3,780,218 2,352,036 1,901,392 1,605,763 1,261,233 1,144,843 1,089,543 1,090,125 1,019,624 971,847 Ens2T 748,854 398,213 285,049 212,246 127,299 96,295 85,680 84,944 69,559 57,513 Ens3T 1,278,533 679,571 485,827 363,881 215,499 165,789 145,762 145,099 117,748 97,375 Ens4T 1,882,653 1,162,076 927,808 783,704 607,545 549,077 526,459 522,067 485,770 462,900 Ens5T 2,611,916 1,723,181 1,441,975 1,250,387 1,033,138 960,309 927,412 923,419 879,444 849,958 Average 2,247,651 1,324,120 1,030,144 840,985 616,494 540,797 506,376 506,967 462,583 431,411
Table 13: M5 estimated costs for each ensemble method and retrain scenario combination (in \$).
The costs tables 14 and 15 show the estimated cost in real values of each scenario for VN1 weekly data.
Method 1 2 3 4 6 8 10 13 26 52 LR 15,234 7,839 5,731 4,292 3,205 2,508 2,140 2,190 1,463 1,099 RF 1,605,768 799,597 562,896 398,954 281,246 199,214 158,959 165,975 81,529 38,865 XGBoost 34,254 17,687 12,796 9,413 6,944 5,262 4,473 4,534 2,863 2,005 LGBM 505,304 260,721 186,904 138,460 101,775 76,738 65,013 66,335 40,721 27,701 CatBoost 52,021 26,594 18,515 13,366 9,611 7,022 5,798 5,977 3,311 2,023 MLP 62,102 34,489 26,431 21,149 17,121 14,391 13,027 13,015 10,255 8,891 LSTM 82,927 49,019 38,851 31,962 27,022 23,465 21,829 21,932 18,415 16,652 TCN 72,763 41,291 31,910 25,464 20,791 17,579 15,983 16,365 12,936 11,284 NBEATSx 80,337 44,837 34,283 27,179 22,057 18,387 16,686 16,702 13,085 11,277 NHITS 80,776 45,458 34,688 27,585 22,336 18,607 16,888 16,855 13,405 11,601 Average 259,149 132,753 95,301 69,782 51,211 38,317 32,080 32,988 19,798 13,140
Table 14: VN1 estimated costs for each base method and retrain scenario combination (in \$).
Method 1 2 3 4 6 8 10 13 26 52 Ens2A 142,439 79,326 60,714 48,328 39,177 32,778 29,713 29,717 23,340 20,167 Ens3A 223,214 124,784 95,402 75,913 61,513 51,385 46,602 46,572 36,745 31,769 Ens4A 1,828,982 924,381 658,299 474,867 342,760 250,599 205,561 212,547 118,274 70,634 Ens5A 1,863,237 942,068 671,095 484,280 349,703 255,861 210,033 217,080 121,137 72,638 Ens2T 49,488 25,526 18,527 13,705 10,149 7,769 6,613 6,724 4,327 3,103 Ens3T 101,509 52,120 37,042 27,071 19,760 14,792 12,410 12,700 7,637 5,126 Ens4T 163,611 86,609 63,473 48,220 36,880 29,183 25,437 25,715 17,892 14,017 Ens5T 236,374 127,900 95,383 73,684 57,671 46,762 41,420 42,080 30,828 25,301 Average 576,106 295,339 212,491 155,758 114,701 86,140 72,223 74,141 45,022 30,344
Table 15: VN1 estimated costs for each ensemble method and retrain scenario combination (in \$).

### v2

@@
 Given the continuous increase in dataset sizes and the complexity of forecasting models, the trade-off between forecast accuracy and computational cost is emerging as an extremely relevant topic, especially in the context of ensemble learning for time series forecasting.
-To asses it, we evaluated ten base models and eight ensemble configurations across two large-scale retail datasets (M5 and VN1), considering both point and probabilistic accuracy under varying retraining frequencies.
+To assess it, we evaluated ten base models and eight ensemble configurations across two large-scale retail datasets, considering both point and probabilistic accuracy under varying retraining frequencies.
 We showed that ensembles consistently improve forecasting performance, particularly in probabilistic settings.
@@
 One of the most intriguing and empirically validated findings in the literature is that simple ensemble strategies, such as taking the mean or median of forecasts, often outperform more sophisticated combination methods like stacking (also known as meta-learning) or weighted ensembles.
-This phenomenon, known as the forecast combination puzzle (Claeskens et al., 2016), suggests that in many cases the added complexity of estimating optimal weights or training a meta-learner does not yield meaningful gains in accuracy and can even lead to overfitting, especially when the number of forecast points is limited.
+This phenomenon, known as the forecast combination puzzle (Claeskens et al., 2016), suggests that in many cases, the added complexity of estimating optimal weights or training a meta-learner does not yield meaningful gains in accuracy and can even lead to overfitting, especially when the number of forecast points is limited.
 Indeed, the success of simple combinations is partly due to their ability to remain robust in the presence of model misspecification, parameter uncertainty, and changing data dynamics, issues that more complex ensemble methods may struggle to accommodate effectively.
-In general, the benefits of ensemble learning are numerous: it improves forecast accuracy, enhances model stability across time and datasets, and increases resilience to structural breaks or shifts in the data.
+In general, the benefits of ensemble learning are numerous: it improves the forecast accuracy, enhances the model stability across time and datasets, and increases resilience to structural breaks or shifts in the data.
 Ensembles also provide a practical hedge against model uncertainty, allowing practitioners to rely less on the assumption that any single model is correctly specified.
@@
 Moreover, in cases where the component models are very similar or highly correlated, the marginal benefit of ensembling diminishes, potentially offering little improvement over the best individual model (Kolassa, 2011).
-Therefore, one of the major advantages of adopting a global modeling approach can be offset when ensemble learning is introduced.
+Therefore, one of the major advantages of adopting a global modeling approach, namely scalability, can be offset when ensemble learning is introduced in the forecasting system.
 Combining several models, even with simple techniques like averaging, inherently increases the computational time, as multiple models must be trained, validated, and maintained, producing substantial computational overhead.
@@
 However, despite these challenges, ensemble learning remains one of the most widely used strategies in the forecasting industry, which further amplifies concerns about the long-term sustainability and scalability of forecasting systems.
-Indeed, from a practical perspective, forecast combinations, coupled with frequent retraining, has also significant environmental costs, since the energy consumption of model training extends beyond direct computational costs, contributing to energy consumption and carbon emissions (Schwartz et al., 2020).
-Therefore, understanding the cost of ensemble forecasting models and the effects of retraining on their performance is of paramount importance for advancing more sustainable forecasting practices.
+Indeed, from a practical perspective, forecast combinations, coupled with frequent retraining, has also significant environmental costs, since the energy consumption of model training extends beyond direct computational costs, contributing to global energy consumption and carbon emissions (Schwartz et al., 2020).
+Therefore, understanding the cost of ensembling forecasting models and the effects of retraining on their performance are of paramount importance for advancing more sustainable forecasting practices.
 1.1 Research Question
@@
 To this end, we use ten global models as base learners, ranging from traditional machine learning methods to advanced deep learning architectures.
-In addition, we evaluate eight ensemble approaches, all trained and tested on the two most recent and comprehensive retail forecasting datasets: the M5 and VN1 competition datasets.
-We also investigated the effects of retraining on the forecasting performance of ensembles, compared to that of the base models.
+In addition, we simulate eight ensemble approaches, all trained and tested on the two most recent and comprehensive retail forecasting datasets: the M5 and VN1 competition datasets.
+We also investigate the effects of retraining on the forecasting performance of ensembles, compared to that of the base models.
 In particular, we assess whether reducing the frequency of retraining, by avoiding re-estimation of base models as new data becomes available, can effectively enhance the cost-accuracy trade-off in ensemble forecasting.
-Hence, we explored a wide range of retraining strategies, from continuous retraining to a no-retraining approach, including various intermediate periodic retraining schemes to encompass the most practical and effective scenarios.
+Hence, we explore a wide range of retraining strategies, from continuous retraining to a no-retraining approach, including various intermediate periodic retraining schemes to encompass the most practical and effective scenarios.
 1.2 Contributions
@@
   • 2.
-    We analyze different ensemble strategies to asses the effects of the ensemble size on forecast accuracy and costs.
+    We analyze different ensemble strategies to assess the effects of the ensemble size on forecast accuracy and costs.
   • 3.
@@
 2 Related works
-The cross-learning approach in time series forecasting has witnessed substantial advancements in recent years, becoming a central theme in contemporary research.
-Today, global models are commonly used as benchmarks in empirical studies, underscoring their growing relevance in the field Semenoglou et al.
-(2021).
+The cross-learning (global) approach in time series forecasting has witnessed substantial advancements in recent years becoming a central theme in contemporary research.
+Today, global models are commonly used as benchmarks in empirical studies, underscoring their growing relevance in the field (Semenoglou et al., 2021).
 Additionally, theoretical contributions from Montero-Manso & Hyndman (2021) and Montero-Manso (2023) have shown that global models can achieve accuracy comparable to local models with reduced complexity and without relying on assumptions about data similarity.
-Indeed, global models have proven particularly effective across a variety of forecasting domains, especially in retail forecasting domain (Makridakis et al., 2022a), and several methods have been proposed to further boost their performance (Godahewa et al., 2021a; Bandara et al., 2020, 2021).
+Indeed, global models have proven particularly effective across a variety of forecasting domains, especially in retail forecasting (Makridakis et al., 2022a), and several methods have been proposed to further boost their performance (Godahewa et al., 2021a; Bandara et al., 2020, 2021).
 In particular, ensemble learning is a powerful tool in the forecasters’ hands offering a reliable strategy to enhance accuracy and robustness by combining the strengths of multiple models, not only in the cross-learning context (Wang et al., 2023).
@@
 However, in applications such as supply chain management, quantifying uncertainty is essential, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022).
-In this regard, the Conformal Inference framework introduced by Vovk et al.
-(2005) offers a flexible tool for uncertainty quantification, and has been successfully applied to time series forecasting problems (Stankeviciute et al., 2021).
 Forecast combinations in the context of probabilistic predictions are an active area of research.
@@
 Other works, such as Spiliotis & Petropoulos (2024), explored retraining strategies and parameter updates for local exponential smoothing models, while Huber & Stuckenschmidt (2020) examined retraining in retail demand forecasting using a limited range of models and datasets.
-Although these studies offer useful insights, direct analysis on the effects of retraining ensemble models remains unexplored.
+Although these studies offer useful insights, direct analysis on the effects of retraining on ensemble models remains unexplored.
 The question of whether ensemble methods are consistently worthwhile is relevant both for the forecasting and the broader machine learning community.
@@
 Our study directly addresses this gap by systematically evaluating different ensemble strategies’ cost-accuracy trade-off.
-We also assess how different updating frequencies affect the forecasting accuracy across the ensemble models, with the aim of testing whether less frequent retraining can be an effective tool to manage this trade-off.
+We also assess how different retraining frequencies affect the forecasting accuracy across the ensemble models, with the aim of testing whether less frequent retraining can be an effective tool to manage this trade-off.
 Answering these questions, this research provides both theoretical clarity and practical guidelines towards more sustainable forecasting practices.
@@
 The M5 competition, part of the M-competitions series led by Spyros Makridakis and collaborators, was designed to benchmark forecasting methods in a retail demand setting (Makridakis et al., 2022b).
-The M5 dataset (Howard & Makridakis, 2020) is widely recognized and extensively studied, comprising 3,049 daily time series representing unit sales of Walmart products.
+The M5 dataset (Howard & Makridakis, 2020) is widely recognized and extensively studied, comprising 30,490 daily time series representing unit sales of Walmart products.
 These sales span three main product categories, Food, Hobbies, and Household, across ten stores in three U.S.
@@
 Unlike the M5 dataset, which contains sales from a single retailer (Walmart) and a limited number of physical stores, the VN1 dataset aggregates sales across 328 warehouses operated by 46 distinct retailers.
-To the best of our knowledge, we are among the first to benchmark forecasting models on this dataset.
+To the best of our knowledge, we are among the first to benchmark ensemble forecasting models on this dataset.
 Together, these two datasets represent the most recent and comprehensive publicly available collections of time series related to retail demand, providing a strong foundation for generalizing our findings in the domain of demand forecasting.
-Table 1: The M5 and the VN1 datasets used in the experiments.
-Dataset Frequency N. Series Min Obs per Series
-M5 Daily (7) 28.298 730
-VN1 Weekly (52) 15.053 157
-In both datasets we concentrate on the most granular level (individual SKUs), because the potential gains from reduced retraining are greatest at lower aggregation levels.
-To ensure a consistent evaluation strategy (Section 3.3), we filtered the series: daily SKUs retained at least two years of data (more than 730 observations), while weekly SKUs required a minimum of three years (more than 157 observations).
+In both datasets, we concentrated on the most disaggregated level (SKU $\times$ Store), because the potential gains from reduced retraining are greatest at lower aggregation levels.
+To ensure a consistent evaluation strategy (see Section 3.3), we filtered the series: daily SKUs retained at least two years of data (more than 730 observations), while weekly SKUs required a minimum of three years (more than 157 observations).
 3.2 Forecasting models
@@
 Indeed, the model parameters $\Theta$ are shared across all series.
-Table 2: Global forecasting models used in the experiment.
-Machine Learning Deep Learning
-Linear Regression (LR) (Godahewa et al., 2021b) MLP (Rosenblatt, 1958)
-Random Forest (RF) (Breiman, 2001) LSTM (Hochreiter & Schmidhuber, 1997)
-XGBoost (Chen & Guestrin, 2016) TCN (Van den Oord et al., 2016)
-LGBM (Ke et al., 2017) NBEATS (Oreshkin et al., 2020)
-CatBoost (Prokhorenkova et al., 2018) NHITS (Challu et al., 2022)
 To ensure a comprehensive evaluation of the ensemble models, we included both traditional machine learning models and state-of-the-art deep learning techniques as base learners, chosen for their proven effectiveness in time series forecasting and their methodological diversity.
-In particular, Table 2 shows the ten different global forecasting models we trained in our experiments.
+Additionally, given the importance of evaluating forecasting models against established benchmarks in time series analysis, we adopted two well-known global models, Linear (Pooled) Regression and Multi-Layer Perceptron, as reference methods for comparison throughout our experiments (Godahewa et al., 2021b).
 Machine learning models are generally easier to train than deep learning approaches but often rely heavily on extensive and careful feature engineering to achieve strong forecasting performance (Januschowski et al., 2022).
@@
 However, they are typically more challenging to train due to their larger number of hyperparameters, which can significantly influence forecast accuracy (Smyl, 2020).
-As in Zanotti (2025), we implemented simplified feature engineering pipelines inspired by the top-performing solutions in the M5 and VN1 competitions.
+As in Zanotti (2025), we implemented simplified feature engineering pipelines inspired by the top-performing solutions of the M5 and VN1 competitions.
 Our feature set included standard time series features such as lags, rolling and expanding means, as well as calendar-related variables like year, month, week, and day of the week.
@@
 For the M5 dataset, we further included external variables like special events.
-Hyperparameters were chosen based on configurations from leading competition entries when available; otherwise, we used the default values recommended by the respective libraries.
+Hyperparameters were chosen based on configurations from leading competition entries when available.
+Otherwise, we used the default values recommended by the respective software libraries.
 3.2.1 Ensemble learning
@@
 In a general formulation, an ensemble forecasting model combines the outputs of multiple models.
-Using the same notation as before, where $\mathcal{Y}$ is the set of all time series and $\Theta^{(j)}$ denotes the parameters of the $j$-th base model $F^{(j)}$, the ensemble prediction for series $Y_{i}$ at horizon $h$ can be defined as:
+Using the same notation introduced before, where $\mathcal{Y}$ is the set of all time series and $\Theta^{(j)}$ denotes the parameters of the $j$-th base model $F^{(j)}$, the ensemble prediction for series $Y_{i}$ at horizon $h$ can be defined as:
 $Y_{i}^{h}=G\left(F^{(1)}(\mathcal{Y},\Theta^{(1)}),F^{(2)}(\mathcal{Y},\Theta^{(2)}),\dots,F^{(J)}(\mathcal{Y},\Theta^{(J)})\right)\mbox{.}$ (2)
@@
 Despite its simplicity, this method is often effective and overall seems to be preferred compared to other combination techniques (Smyl & Hua, 2019).
-Moreover, equally weighting quantiles through a simple average yields robust and improved forecast skill because the error in estimating optimal weights usually exacerbates the ensemble predictions (Busetti, 2017).
-Lastly, simple combination methods have the lowest computational time possible 111 Simple combinations, such as the mean or median, require only basic arithmetic operations on the forecast outputs of the base models, avoiding the need for model training or optimization, as is required in more complex approaches like meta-learning.
-, allowing us to effectively study, without loss of generality, the associated costs of ensembling.
-Table 3: Composition of ENSACC and ENSTIME ensembles for the M5 and VN1 datasets.
-Dataset Ensemble Models Included
-M5 ENSACC (Accuracy-driven combinations)
-ENS2A XGBoost, LGBM
-ENS3A XGBoost, LGBM, LR
-ENS4A XGBoost, LGBM, LR, NBEATSx
-ENS5A XGBoost, LGBM, LR, NBEATSx, MLP
-ENSTIME (Time-efficient combinations)
-ENS2T CatBoost, LR
-ENS3T CatBoost, LR, XGBoost
-ENS4T CatBoost, LR, XGBoost, MLP
-ENS5T CatBoost, LR, XGBoost, MLP, NBEATSx
-VN1 ENSACC (Accuracy-driven combinations)
-ENS2A MLP, NBEATSx
-ENS3A MLP, NBEATSx, NHITS
-ENS4A MLP, NBEATSx, NHITS, RF
-ENS5A MLP, NBEATSx, NHITS, RF, XGBoost
-ENSTIME (Time-efficient combinations)
-ENS2T LR, XGBoost
-ENS3T LR, XGBoost, CatBoost
-ENS4T LR, XGBoost, CatBoost, MLP
-ENS5T LR, XGBoost, CatBoost, MLP, TCN
-In our study, we employed simple mean ensembles to combine the forecasts of multiple global models, following two distinct selection criteria.
+Moreover, equally weighting quantiles through a simple average yields robust and improved forecast results because the error in estimating optimal weights usually exacerbates the ensemble predictions (Busetti, 2017).
+Lastly, simple combination methods have the lowest computational time possible 111 Simple combinations, such as the mean or median, require only basic arithmetic operations on the forecast outputs of the base models, avoiding the need for model training or optimization, as it is required by more complex approaches like meta-learning.
+, allowing us to effectively study the associated costs of ensembling without any loss of generality.
+In our study, we simulate various ensemble strategies to assess, a posteriori, how forecast performance over the full test window would have changed if forecasting combinations had been used in place of the individual base models.
+Hence, we employed simple mean ensembles to combine the forecasts of multiple pools of global models, following two distinct selection criteria.
 The first strategy, which we refer to as ENSACC, combines the base models that demonstrated the highest individual point forecast accuracy (measured by Equation 4).
@@
 For both strategies, we constructed ensembles of increasing size, combining the top 2, 3, 4, and 5 models according to each criterion.
-Table 3 shows the composition of the eight different ensembles for each dataset.
+This idea of only using a subset of the base learners in ensembling is essentially a variant of model pooling, as defined by Kourentzes et al.
+(2019).
 This tiered design allows us to assess how forecast accuracy and computational efficiency evolve as additional models are added to the ensemble.
@@
 These scenarios are tailored to the frequency of each dataset, which in turn determines the forecast horizon and business review cycles.
-Table 4 summarizes the selected retraining windows, test periods, and forecast horizons.
-Table 4: The retraining scenarios, the test window, and the horizon for the M5 and the VN1 datasets.
-Dataset Frequency Retraining Scenarios (r) Test Window (T) Horizon (h)
-M5 Daily (7) 7, 14, 21, 30, 60, 90, 120, 150, 180, 364 364 28
-VN1 Weekly (52) 1, 2, 3, 4, 6, 8, 10, 13, 26, 52 52 13
+Table 1 summarizes the selected retraining sets, training and test sizes, and forecast horizons.
+Table 1: Retraining set, train and test size, frequency, forecast horizon, and step size for each dataset. The setup is based on the respective time series frequency.
+M5 VN1
+Frequency ($f$) daily (7) weekly (52)
+Train size ($n$) $\geq 730$ $\geq 157$
+Test size ($T$) 364 52
+Horizon ($h$) 28 13
+Step size ($p$) 1 1
+Retrain set ($r$) {7, 14, 21, 30, 60, 90, 120, 150, 180, 364} {1, 2, 3, 4, 6, 8, 10, 13, 26, 52}
 For example, $r=7$ for daily data implies weekly retraining.
@@
 However, for daily data, we consider $r=7$ (weekly retraining) as the practical benchmark, since daily updates are uncommon in real-world settings.
-Conversely, the no-retraining scenario $r=T$, where $T$ is the length of the test set, fits the model only once and uses it for the entire forecasting horizon, minimizing computational load but likely yielding the lowest accuracy.
+Conversely, the no-retraining scenario $r=T$, where $T$ is the length of the test set, fits the model only once, and uses it for the entire forecasting horizon, minimizing computational load but likely yielding the lowest accuracy.
 All intermediate values $1<r<T$ represent periodic retraining strategies, where both forecast accuracy and computational cost are expected to decrease as $r$ increases.
-Since we use global models, the training set at each retraining step includes all time series in the dataset, each extended by $r$ new observations.
-We evaluate two update strategies: (i) full retraining at each window $r$, or (ii) no update between retrain points, i.e., the model remains unchanged for the next $r$ periods.
 3.3.2 Rolling window evaluation
@@
 At each step, the model is trained on the current training set and used to predict the next $h$ observations.
-The forecast origin is then shifted forward by a fixed step, and the process is repeated, either with a growing (expanding) or fixed-size training window.
+The forecast origin is then shifted forward by a fixed step (the step size $p$), and the process is repeated, either with a growing (expanding) or fixed-size training window.
 Performance metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
 Compared to fixed origin evaluation, which offers only a single evaluation point, rolling origin evaluation provides a more robust assessment by capturing performance across varying conditions such as seasonal shifts, level changes, or trend evolutions (Bergmeir & Benítez, 2012).
-This is particularly beneficial in dynamic settings like retail or supply chain management, where models must adapt to frequent changes.
+This is particularly beneficial in dynamic settings like retail or supply chain management, where models need to adapt to frequent changes.
 The setup is flexible: the training window can either expand to include all available past data or remain fixed to a specified length.
 Most practitioners favor the expanding window, especially when time series are short, as it leverages the full data history (Petropoulos & et al., 2022).
-In our study, we adopt the expanding window strategy to better reflect realistic business forecasting practices and to accommodate short series, such as those in the weekly VN1 dataset.
-Following Zanotti (2025), Table 4 outlines the key parameters used in our experiments.
-Lastly, we set the step size to one in all cases to maximize the number of evaluation points across each retraining scenario.
+In our study, we adopt the expanding window strategy to reflect realistic business forecasting practices better and to accommodate short series, such as those in the weekly VN1 dataset.
+Table 1 outlines the key parameters used in our experiments.
+In particular, we used a test set of one complete year to avoid results being affected by intra-year fluctuations, and we set the forecast horizon based on the type of business decisions that usually daily and weekly forecasts support.
+Lastly, we adopted a step size of one to maximize the number of evaluations across each retraining scenario.
 3.4 Evaluation metrics
@@
 Furthermore, due to scale variations across series, employing a scale-independent metric is essential.
-To address these concerns, following Zanotti (2025), we adopted the Root Mean Squared Scaled Error (RMSSE), introduced by Hyndman & Koehler (2006), as our primary point forecast accuracy measure.
+To address these concerns, we adopted the Root Mean Squared Scaled Error (RMSSE), introduced by Hyndman & Koehler (2006), as our primary point forecast accuracy measure.
 RMSSE is calculated as:
@@
 This metric compares the mean squared error of a forecast to that of a seasonal naïve benchmark, providing a relative performance measure.
-It was the official evaluation metric used in the M5 competition (with $s=1$) (Makridakis et al., 2022b), one of the datasets analyzed in our study.
+It was the official evaluation metric used in the M5 competition (with $s=1$, hence using the naive model as a benchmark) (Makridakis et al., 2022b).
 Lower RMSSE values indicate more accurate forecasts.
 Beyond point accuracy, we also assessed the probabilistic quality of forecasts across different retraining scenarios.
-Since most of the machine learning and deep learning models we use do not natively produce probabilistic outputs, we employed Conformal Inference to generate prediction intervals.
-Conformal Inference is a versatile framework that quantifies predictive uncertainty using a calibration (validation) set, without relying on strong distributional assumptions (Vovk et al., 2005).
-Originally developed for i.i.d.
-data, it has recently been adapted to time series contexts (Stankeviciute et al., 2021).
-Its properties, distribution-free, model-agnostic, efficient, and suitable for small datasets, make it especially well-suited for forecasting benchmarks involving diverse models and datasets.
-As in Zanotti (2025), to evaluate the quality of the prediction intervals, we used the Quantile Loss (QL), also known as Pinball Loss, and its average form, the Multi-Quantile Loss (MQL):
-$\text{QL}=\frac{1}{h}\sum_{t=n+1}^{n+h}\left(q\cdot(y_{t}-\hat{y}_{t})\cdot\mathbb{I}_{y_{t}\geq\hat{y}_{t}}+(1-q)\cdot(\hat{y}_{t}-y_{t})\cdot\mathbb{I}_{y_{t}<\hat{y}_{t}}\right)\mbox{,}$ (5)
-$\text{MQL}=\frac{1}{\mathcal{Q}}\sum_{q\in\mathcal{Q}}\text{QL}(q)\mbox{.}$ (6)
+To evaluate the quality of the estimated quantiles, we used a scaled version of the Quantile Loss (SQL), also known as Pinball Loss, and its average form, the scaled Multi-Quantile Loss (SMQL):
+$\text{SQL}=\frac{1}{h}\frac{\sum_{t=n+1}^{n+h}\left(q\cdot(y_{t}-\hat{y}_{t})\cdot\mathbb{I}_{y_{t}\geq\hat{y}_{t}}+(1-q)\cdot(\hat{y}_{t}-y_{t})\cdot\mathbb{I}_{y_{t}<\hat{y}_{t}}\right)}{\frac{1}{n-s}\sum_{t=s+1}^{n}|y_{t}-y_{t-s}|}\mbox{,}$ (5)
+$\text{SMQL}=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\text{SQL}(q)\mbox{,}$ (6)
 As a proper scoring rule, QL enables a rigorous evaluation of probabilistic forecasts (Kolassa, 2016).
-MQL, particularly in its weighted form, was the principal metric in the M5 Uncertainty competition (Makridakis et al., 2022c).
-For our analysis, we constructed prediction intervals around the median and six central levels: 60%, 70%, 80%, 90%, 95%, and 99%, yielding a total of 13 quantiles.
-The lower quantiles (e.g., median, 60%, 70%) describe the forecast center, while the upper quantiles (90% and above) are crucial for assessing tail risks, key to managing safety stock in retail settings (Barrow & Kourentzes, 2016).
+The scaling factor is the in-sample, one-step-ahead mean absolute error of the seasonal Naive model.
+The scaled MQL, particularly in its weighted form and with $s=1$, was the principal metric of the M5 Uncertainty competition (Makridakis et al., 2022c).
+For daily data we used $s=7$, because $s=1$ is arguably an irrelevant benchmark in daily retail data, which typically exhibits strong day-of-week seasonality (Fildes et al., 2022).
+Instead, we set $s=1$ for weekly data.
+For our analysis, we considered a total of 14 different quantiles, $q$, of probability levels {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}.
+Therefore, we studied the 50% 60%, 70%, 80%, 90%, 95%, and 99% symmetric prediction intervals.
+The lower intervals (e.g., 50%, and 60%) describe the forecast center, while the upper intervals (90% and above) are crucial for assessing tail risks, key to managing safety stock in retail settings (Barrow & Kourentzes, 2016).
 These quantiles offer a well-rounded view of forecast uncertainty.
-To ensure valid estimation of predictive uncertainty, conformal prediction intervals were computed on validation sets at least twice the length of the forecast horizon.
-This requirement limited the number of time series we could retain in the dataset, as explained in previous sections.
 In addition to accuracy, we explicitly evaluated the computational cost of generating forecasts by measuring Computing Time (CT), defined as the number of seconds needed to train a model and produce $h$ step-ahead forecasts (Zanotti, 2025).
-CT directly reflects forecasting costs in real-world deployments, especially when using cloud-based infrastructure with pay-as-you-go pricing (Spiliotis & Petropoulos, 2024).
+CT directly reflects forecasting costs in real-world applications, especially when using cloud-based infrastructure with pay-as-you-go pricing (Spiliotis & Petropoulos, 2024).
 Thus, CT was recorded for each model across all retraining scenarios, with lower CT values indicating more efficient forecasting.
-To facilitate consistent comparisons, results for all evaluation metrics were normalized relative to the baseline (most frequent) retraining scenario for each dataset frequency.
+To facilitate consistent comparisons, results for all evaluation metrics were normalized relative to the baseline (continuous) retraining scenario for each dataset frequency.
 To statistically assess performance differences across scenarios, we applied the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).
@@
 We draw insights from both the M5 and VN1 datasets to identify consistent patterns and dataset-specific nuances.
-Table 5: Overall forecasting performance and computational cost for each method across datasets. RMSSE, MQL, and CT (in seconds). Minimum values per column are highlighted in bold.
+Table 2: Updated forecasting performance and computational cost for each method across datasets.
+RMSSE, SMQL, and CT (in seconds).
+Minimum RMSSE and SMQL values per column are highlighted in bold.
+The baseline scenarios are $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
 Type Method M5 VN1
-RMSSE MQL CT RMSSE MQL CT
-ML LR 0.777 0.267 11373 6.549 2.896 236
-RF 1.868 2.590 24862
-XGBoost 0.755 0.258 15417 1.890 2.469 530
-LGBM 0.771 0.256 44429 3.542 2.625 7824
-CatBoost 0.947 0.263 10424 5.762 2.845 805
-DL MLP 0.821 0.281 17584 1.543 2.492 962
-LSTM 1.913 2.843 1284
-TCN 0.865 0.290 33364 1.913 2.843 1127
-NBEATSx 0.815 0.279 21226 1.698 2.626 1244
-NHITS 0.828 0.284 21969 1.699 2.632 1251
-ENSACC Ens2A 0.757 0.255 59846 1.472 2.369 2205
-Ens3A 0.757 0.256 71219 1.517 2.410 3456
-Ens4A 0.758 0.249 92445 1.524 2.386 28318
-Ens5A 0.763 0.251 110029 1.544 2.375 28849
-ENSTIME Ens2T 0.814 0.230 21797 5.719 2.545 766
-Ens3T 0.776 0.223 37214 5.536 2.588 1572
-Ens4T 0.764 0.218 54797 4.472 2.419 2533
-Ens5T 0.763 0.220 76024 3.672 2.339 3660
+RMSSE SMQL CT RMSSE SMQL CT
+ML LR 0.760 0.269 11373 6.535 4.005 236
+RF 1.871 1.220 24862
+XGBoost 0.739 0.246 15417 1.898 1.503 530
+LGBM 0.755 0.247 44429 3.540 2.318 7824
+CatBoost 0.926 0.280 10424 5.355 3.889 805
+DL MLP 0.804 0.264 17584 1.554 0.806 962
+LSTM 1.960 0.892 1284
+TCN 0.847 0.271 33364 1.960 0.891 1127
+NBEATSx 0.799 0.263 21226 1.704 0.991 1244
+NHITS 0.812 0.267 21969 1.705 0.992 1251
+ENSACC Ens2A 0.741 0.244 59846 1.478 0.826 2205
+Ens3A 0.741 0.248 71219 1.523 0.867 3456
+Ens4A 0.743 0.241 92445 1.530 0.926 28318
+Ens5A 0.748 0.242 110029 1.550 1.013 28849
+ENSTIME Ens2T 0.798 0.257 21797 5.694 4.388 766
+Ens3T 0.762 0.249 37214 5.286 3.972 1572
+Ens4T 0.751 0.242 54797 4.447 3.420 2533
+Ens5T 0.750 0.243 76024 3.664 2.832 3660
 Table 2 summarizes the performance and computational time of all forecasting methods evaluated in this study, across both the M5 and VN1 datasets.
 The models are grouped into four categories: Machine Learning (ML), Deep Learning (DL), ENSACC (ensembles optimized for accuracy), and ENSTIME (ensembles optimized for computational efficiency).
-For each method, the table reports three key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to assess point forecast accuracy, MQL (Multi-Quantile Loss) to evaluate probabilistic forecast performance, and CT (Computing Time, in seconds) to reflect the computational time under a cloud-based setting.
+For each method, the table reports three key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to assess point forecast accuracy, SMQL (Scaled Multi-Quantile Loss) to evaluate probabilistic forecast performance, and CT (Computing Time, in seconds) to reflect the computational time under a cloud-based setting.
 Overall 222 The overall results are derived from the baseline scenario, that is $r=7$ for M5 and $r=1$ for VN1, as this setting is regarded as standard in both theoretical and practical applications.
@@
 Interestingly, ENSTIME ensembles show increasing returns in accuracy as more models are added: their performance consistently improves with each additional component.
-Indeed, in contrast to ENSACC ensembles, where combining just two models is often sufficient to achieve optimal performance (as illustrated in Figure 1), ENSTIME combinations benefit from larger ensemble sizes, with forecast accuracy continuing to improve as more low-cost models are included.
-Figure 1: Friedman-Nemenyi test results of ensembles based on RMSSE.
-A similar pattern emerges for probabilistic accuracy evaluated using Multi-Quantile Loss (MQL).
+Indeed, in contrast to ENSACC ensembles, where combining just two models is often sufficient to achieve optimal performance, ENSTIME combinations benefit from larger ensemble sizes, with forecast accuracy continuing to improve as more low-cost models are included (see Supplementary material for significance tests).
+A similar pattern emerges for probabilistic accuracy evaluated using Scaled Multi-Quantile Loss (SMQL).
 Compared to base models, ensembles achieve superior uncertainty quantification, benefiting from the diversity in quantile predictions among different models.
@@
 On the contrary, increasing the number of models in the ensemble based on computational efficiency leads to consistently improved probabilistic forecasting performance, which may even outperform the ENSACC combinations.
-As shown in Figure 2, ENSTIME ensembles achieve significantly better probabilistic forecasting performance on the M5 dataset compared to all accuracy-based combinations.
+ENSTIME ensembles achieve similar probabilistic forecasting performance on the M5 dataset compared to all accuracy-based combinations, while their performance is worse on the VN1 dataset (see also Supplementary material).
 In contrast, for ENSACC ensembles, combining more than two base models does not yield additional improvements, confirming once again that a small, high-performing subset is sufficient in that case.
-Figure 2: Friedman-Nemenyi test results of ensembles based on MQL.
 Unsurprisingly, ensemble models require significantly more computing time than individual models, with algorithms like LGBM and RF contributing disproportionately to this overhead.
@@
 For most ensemble models, particularly under low-frequency retraining (i.e., periodic updates), performance is nearly indistinguishable from the baseline.
-Although slight accuracy degradation is observed for the ENSTIME ensembles under less frequent retraining, it remains modest—below 2% in M5 and under 5% in VN1, even in the no-retraining condition.
-These findings build upon and extend the results of Zanotti (2025), suggesting that reducing retraining frequency does not significantly harm the point forecast accuracy of global model ensembles.
+Although slight accuracy degradation is observed for the ENSTIME ensembles under less frequent retraining, it remains modest, below 2% in M5 and under 3% in VN1, even in the no-retraining condition.
+These findings build upon and extend the results of Zanotti (2025), suggesting that reducing retraining frequency does not significantly harm the point forecast accuracy of global ensemble models.
 This robustness likely stems from the relative stability of the datasets, which do not exhibit substantial structural changes or concept drift.
@@
 Figure 1: RMSSE results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
-Similarly, Figure 4 illustrates the relative accuracy of the ensemble models in a probabilistic forecasting context, as measured by the Multi-Quantile Loss (MQL).
+Similarly, Figure 2 illustrates the relative accuracy of the ensemble models in a probabilistic forecasting context, as measured by the Scaled Multi-Quantile Loss (SMQL).
 For the M5 dataset, we observe a clear trend where probabilistic accuracy decreases as retraining becomes less frequent, indicating that regular updates are beneficial for maintaining high-quality uncertainty estimates.
 This pattern holds across the different ensemble strategies.
-However, for ENSACC ensembles, the decline in accuracy is minimal for frequent retraining scenarios and becomes slightly more noticeable at higher levels of retraining, though it remains within 5 percentage points.
-A consistent performance gap is also evident between ENSACC and ENSTIME ensembles, with the former outperforming the latter as the retraining scenario increases.
-This difference is largely attributable to the inclusion of CatBoost in the ENSTIME ensembles, a model that shows relatively weak performance in this setting.
+For ENSACC ensembles the decline in accuracy is minimal for frequent retraining scenarios and becomes slightly more noticeable at higher levels of retraining, though it remains within 4 percentage points.
+ENSTIME combinations, instead, show even higher performance, in particular in the case of the smallest ensembles.
 In contrast, the VN1 dataset reveals a different behavior.
 Here, the relationship between retraining frequency and probabilistic accuracy follows a near-convex pattern.
-Performance initially improves with less frequent retraining, peaking around $r=4$, and begins to decline thereafter.
+Performance initially improves with less frequent retraining, peaking around $r=8$, and begins to decline thereafter.
 This suggests that very frequent updates may not be necessary for probabilistic performance in stable, lower-frequency datasets like VN1.
-Nevertheless, ENSACC ensembles continue to outperform ENSTIME ones across the retraining scenarios.
 Interestingly, the rate of performance deterioration with increased retraining window is comparable across both ensemble strategies in VN1, indicating that the marginal impact of retraining frequency is relatively uniform across the proposed ensemble techniques.
 Also these results on the probabilistic performance of ensembles are comparable to those obtained by Zanotti (2025) on the base models.
-Figure 4: MQL results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
-The Friedman-Nemenyi tests, used to compare the accuracy produced by the same model over different retraining scenarios, confirm the previous results (see Supplementary material).
+Figure 2: SMQL results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
+The Friedman-Nemenyi tests, used to compare the accuracy produced by the same model over different retraining scenarios, confirm the above results (see Supplementary material).
 Specifically, in the evaluation of both point forecast accuracy and probabilistic forecast accuracy, some level of periodic retraining is at least as good as the continuous retraining scenario.
@@
 These findings have direct implications for forecasting cost management, emphasizing the importance of considering retraining frequency and ensemble size when selecting models for large-scale operational use.
-Overall, the evidence from Figures 3, 4, and 5, along with the statistical significance tests, indicates that reducing the retraining frequency of ensemble models does not negatively impact their predictive accuracy.
+Overall, the evidence from Figures 1, 2, and 3, along with the statistical significance tests, indicates that reducing the retraining frequency of ensemble models does not negatively impact their predictive accuracy.
 At the same time, it lowers the computational time required to generate forecasts.
@@
 As previously noted, computational time can be directly translated into monetary cost for organizations.
-In line with Nikolopoulos & Petropoulos (2018); Fotios Petropoulos & Spiliotis (2024); Zanotti (2025), we adopted standard pricing assumptions for cloud computing services to estimate both the overall cost of forecasting and the specific costs associated with each retraining scenario for the base global models and each ensemble 333 It is important to note that the costs have been normalized by the number of SKUs in each dataset, allowing for direct comparison across datasets.
-This normalization enables meaningful conclusions to be drawn regarding the impact of time series frequency on forecasting costs.
+In line with Nikolopoulos & Petropoulos (2018); Fotios Petropoulos & Spiliotis (2024); Zanotti (2025), we adopted standard pricing assumptions for cloud computing services to estimate both the overall cost of forecasting and the specific costs associated with each retraining scenario for the base global models and each ensemble 333 The costs have been normalized by the number of SKUs in each dataset to allow comparisons.
 .
-Figures 6, 7, 8 9, and 10 illustrate the estimated forecasting costs for a large-scale retailer, assuming a computing service rate of \$3.5 per hour, with 200,000 unique SKUs and 5,000 stores.
-Total forecast production cost, computed by combining training time and forecast generation under a cloud-computing pricing model, highlights the trade-off central to our study.
+Figures 4, 5, 6, and 7 illustrate the estimated forecasting costs for a large-scale retailer, assuming a computing service rate of \$3.5 per hour, with 200,000 unique SKUs and 5,000 stores.
+The total forecast production cost (Figure 4), computed by combining training time and forecast generation under a cloud-computing pricing model, highlights the trade-off central to our study.
 ENSACC ensembles incur the highest costs, rapidly reaching millions of dollars under frequent retraining.
@@
 Moreover, as dataset size increases, the trade-off between forecast accuracy and cost becomes significantly more pronounced.
-This is clearly illustrated in Figures 7 and 8.
-For the VN1 dataset, ensemble methods offer noticeable improvements in both point and probabilistic forecasting accuracy compared to most individual base models, while incurring only a modest increase in cost.
+This is clearly illustrated in Figures 5 and 6.
+For the VN1 dataset, ENSACC methods offer noticeable improvements in both point and probabilistic forecasting accuracy compared to most individual base models, while incurring only a modest increase in cost.
 In this context, there is little reason to favor ENSTIME ensembles over ENSACC ones, as the cost difference is minor and the performance gain from accuracy-based combinations is evident.
@@
 Here, ensemble models, particularly those optimized for accuracy, are significantly more expensive than individual models.
-Additionally, the performance advantage of ENSACC ensembles over the most accurate base models is minimal in terms of both RMSSE and MQL.
+Additionally, the performance advantage of ENSACC ensembles over the most accurate base models is minimal in terms of both RMSSE and SMQL.
 In contrast, ENSTIME ensembles can achieve comparable levels of accuracy with substantially lower additional cost, making them a more attractive option.
@@
 Notably, forecast combinations that disregard the computational efficiency of their base models can lead to excessive and unjustified costs.
-This is particularly evident in the case of Ens4A and Ens5A on the VN1 dataset, where the inclusion of Random Forest (a highly expensive model in this context) results in disproportionately high forecasting costs with only marginal accuracy improvements.
+This is particularly evident in the case of Ens4A and Ens5A.
+For instance, on the VN1 dataset, where the inclusion of Random Forest (a highly expensive model in this context) results in disproportionately high forecasting costs with only marginal accuracy improvements.
 Figure 4: Cost analysis. Overall cost results for the M5 and VN1 datasets.
 Figure 5: Cost analysis. RMSSE vs Cost (\$) for the M5 and VN1 datasets.
-Figure 8: Cost analysis. MQL vs Cost (\$) for the M5 and VN1 datasets.
-Finally, 9 and 10 show the comparison between base models and ensembles across all retraining scenarios.
+Figure 6: Cost analysis. SMQL vs Cost (\$) for the M5 and VN1 datasets.
+Finally, 7 shows the comparison between base models and ensembles across all retraining scenarios.
 As expected, forecasting costs decrease exponentially with less frequent retraining.
@@
 Moreover, the cost-saving benefits of reduced retraining frequency become more pronounced as dataset size increases.
-In the VN1 dataset, cost differences between retraining scenarios remain relatively small, even for ensemble models, suggesting limited sensitivity to retraining frequency.
+In the VN1 dataset, cost differences between retraining scenarios remain relatively small, even for ensemble models, suggesting limited sensitivity to retraining frequency (see Supplementary material).
 In contrast, the M5 dataset exhibits steady, incremental reductions in cost as retraining becomes less frequent.
@@
 Each method and retrain scenario combination are shown in relative terms with respect to the baseline scenario, $r=7$.
-Figure 10: Cost analysis.
-Comparison between base models and ensembles for the VN1 dataset.
-Each method and retrain scenario combination are shown in relative terms with respect to the baseline scenario, $r=1$.
-One might argue that, for a large retailer, these costs (and the potential savings) are relatively minor.
+One might argue that, for a large retailer, these costs (and the potential savings) are relatively small.
 However, it’s important to emphasize that the cost reductions achieved through a combination of less frequent retraining, smaller ensemble sizes, and more efficient model combinations typically come with no loss in forecasting accuracy, particularly for point forecasts.
@@
 This comprehensive analysis reinforces the core message of the article: ensemble methods offer strong accuracy benefits, but their cost can be substantial, especially under frequent retraining.
-Balancing accuracy and sustainability requires a deliberate model selection and retraining design, with lightweight ensembles, like ENSTIME, being a possible solution in large-scale forecasting systems.
+Balancing accuracy and sustainability requires deliberate model selection and retraining design, with lightweight ensembles, like ENSTIME, being a possible solution in large-scale forecasting systems.
 5 Conclusions
@@
 Supplementary material
-In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets.
-The Tables 6 and 7 show the forecast accuracy of the different ensembles along the examined retrain scenarios for the M5 daily dataset, while Table 8 depicts the computing time in seconds.
+In this section, we provide supplementary details on the design of our experiment, together with tables and figures related to the empirical results of the M5 and VN1 datasets.
+Conformal inference
+The Conformal Inference framework introduced by Vovk et al.
+(2005) offers a flexible tool for uncertainty quantification, and has been successfully applied to time series forecasting problems (Stankeviciute et al., 2021).
+Since most of the machine learning and deep learning models we use do not natively produce probabilistic outputs, we employed Conformal Inference to generate prediction intervals.
+Conformal Inference is a versatile framework that quantifies predictive uncertainty using a calibration (validation) set, without relying on strong distributional assumptions (Vovk et al., 2005).
+Originally developed for i.i.d.
+data, it has recently been adapted also to time series contexts (Stankeviciute et al., 2021).
+Its properties, distribution-free, model-agnostic, efficient, and suitable for small datasets, make it especially well-suited for forecasting benchmarks involving diverse models and datasets.
+An alternative approach is to train models using quantile loss functions to directly produce quantile forecasts.
+While this method can effectively capture uncertainty, as demonstrated in the M5 Uncertainty competition (Makridakis et al., 2022c; Lainder & Wolfinger, 2022), it requires training a separate model for each quantile level.
+This significantly increases the computational burden, making it less suitable for large-scale forecasting applications where efficiency and sustainability are critical considerations.
+Moreover, although multi-quantile loss functions can generate multiple quantiles and the corresponding prediction intervals, the resulting intervals are often uncalibrated, that is, the proportion of observations falling within them may not match the intended confidence level.
+For these multiple reasons, we chose to use Conformal Inference as a general framework for producing quantile forecasts.
+To ensure valid estimation of predictive uncertainty, conformal prediction intervals were computed on validation sets at least twice the length of the forecast horizon.
+This requirement limited the number of time series we could retain in the dataset as explained in previous sections We used a validation set four times larger than the forecast horizon for the M5 dataset and double the forecast horizon for the VN1 dataset.
+This limited the number of series from 30,490 to 28,298 for the M5 dataset.
+Empirical results
+Table 3 shows the ten different global forecasting models we trained in our experiments.
+For the cost of computation, the Random Forest and the LSTM models have been tested on the VN1 weekly dataset only.
+Table 4 shows the composition of the eight different ensembles for each dataset.
+The Tables 5 and 6 show the forecast accuracy of the different ensembles along the examined retrain scenarios for the M5 daily dataset, while Table 7 depicts the computing time in seconds.
+The Tables 8 and 9 show the forecast accuracy of the different ensembles along the examined retrain scenarios for the VN1 weekly dataset, while Table 10 depicts the computing time in seconds.
+Figures 8 and 9 show the results of the Friedman-Nemenyi test for both point forecast and probabilistic accuracy for the M5 and VN1 datasets, with respect to the baseline scenario.
+Figures 10 and 11 show the results of the Friedman-Nemenyi test in the context of point forecast accuracy for the M5 and VN1 datasets respectively.
+Figures 12 and 13 present the test results related to the probabilistic evaluation.
+The costs tables 11 and 12 show the estimated cost in real values of each scenario for M5 daily data.
+The costs tables 13 and 14 show the estimated cost in real values of each scenario for the VN1 weekly data.
+Figure 14 shows the cost profiles of each model across the retraining scenarios for the VN1 dataset.
+Table 3: Global forecasting models used as base models in the experiment.
+Machine Learning Deep Learning
+Linear Regression (LR) (Godahewa et al., 2021b) MLP (Rosenblatt, 1958)
+Random Forest (RF) (Breiman, 2001) LSTM (Hochreiter & Schmidhuber, 1997)
+XGBoost (Chen & Guestrin, 2016) TCN (Van den Oord et al., 2016)
+LGBM (Ke et al., 2017) NBEATS (Oreshkin et al., 2020)
+CatBoost (Prokhorenkova et al., 2018) NHITS (Challu et al., 2022)
+Table 4: Composition of ENSACC and ENSTIME ensembles for the M5 and VN1 datasets.
+Dataset Ensemble Models Included
+M5 ENSACC (Accuracy-driven combinations)
+ENS2A XGBoost, LGBM
+ENS3A XGBoost, LGBM, LR
+ENS4A XGBoost, LGBM, LR, NBEATSx
+ENS5A XGBoost, LGBM, LR, NBEATSx, MLP
+ENSTIME (Time-efficient combinations)
+ENS2T CatBoost, LR
+ENS3T CatBoost, LR, XGBoost
+ENS4T CatBoost, LR, XGBoost, MLP
+ENS5T CatBoost, LR, XGBoost, MLP, NBEATSx
+VN1 ENSACC (Accuracy-driven combinations)
+ENS2A MLP, NBEATSx
+ENS3A MLP, NBEATSx, NHITS
+ENS4A MLP, NBEATSx, NHITS, RF
+ENS5A MLP, NBEATSx, NHITS, RF, XGBoost
+ENSTIME (Time-efficient combinations)
+ENS2T LR, XGBoost
+ENS3T LR, XGBoost, CatBoost
+ENS4T LR, XGBoost, CatBoost, MLP
+ENS5T LR, XGBoost, CatBoost, MLP, TCN
 Method 7 14 21 30 60 90 120 150 180 364
-Ens2A 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757
-Ens3A 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757 0.757
-Ens4A 0.758 0.758 0.758 0.759 0.758 0.758 0.758 0.758 0.759 0.759
-Ens5A 0.763 0.763 0.763 0.764 0.763 0.764 0.763 0.764 0.764 0.764
-Ens2T 0.814 0.827 0.814 0.826 0.822 0.830 0.826 0.827 0.828 0.824
-Ens3T 0.776 0.786 0.776 0.786 0.784 0.787 0.785 0.786 0.787 0.785
-Ens4T 0.764 0.773 0.764 0.773 0.772 0.774 0.772 0.773 0.773 0.772
-Ens5T 0.763 0.771 0.762 0.772 0.771 0.772 0.770 0.771 0.772 0.771
+Ens2A 0.741 0.741 0.741 0.741 0.741 0.740 0.740 0.740 0.740 0.740
+Ens3A 0.741 0.741 0.741 0.741 0.741 0.740 0.741 0.741 0.740 0.740
+Ens4A 0.743 0.742 0.742 0.743 0.742 0.742 0.742 0.742 0.742 0.742
+Ens5A 0.748 0.747 0.747 0.748 0.747 0.747 0.747 0.747 0.747 0.747
+Ens2T 0.798 0.810 0.799 0.809 0.805 0.812 0.809 0.809 0.810 0.806
+Ens3T 0.762 0.770 0.762 0.769 0.768 0.771 0.769 0.769 0.770 0.767
+Ens4T 0.751 0.756 0.751 0.756 0.756 0.757 0.755 0.756 0.756 0.755
+Ens5T 0.750 0.754 0.749 0.756 0.754 0.755 0.753 0.754 0.754 0.753
 Table 5: M5 RMSSE values for each method and retrain scenario combination.
 Method 7 14 21 30 60 90 120 150 180 364
-Ens2A 0.255 0.255 0.256 0.256 0.258 0.258 0.259 0.259 0.259 0.262
-Ens3A 0.256 0.256 0.257 0.257 0.259 0.259 0.260 0.260 0.261 0.263
-Ens4A 0.249 0.250 0.251 0.251 0.253 0.254 0.255 0.255 0.256 0.259
-Ens5A 0.251 0.251 0.252 0.253 0.255 0.256 0.257 0.257 0.259 0.261
-Ens2T 0.230 0.265 0.228 0.267 0.267 0.269 0.268 0.269 0.269 0.270
-Ens3T 0.223 0.257 0.221 0.258 0.260 0.261 0.261 0.261 0.262 0.263
-Ens4T 0.218 0.253 0.216 0.255 0.256 0.257 0.257 0.258 0.259 0.261
-Ens5T 0.220 0.254 0.218 0.257 0.258 0.259 0.260 0.260 0.262 0.264
-Table 7: M5 MQL values for each method and retrain scenario combination.
+Ens2A 0.244 0.244 0.245 0.245 0.246 0.247 0.248 0.247 0.248 0.249
+Ens3A 0.248 0.249 0.249 0.249 0.251 0.251 0.253 0.252 0.253 0.254
+Ens4A 0.241 0.242 0.242 0.243 0.245 0.246 0.247 0.246 0.248 0.250
+Ens5A 0.242 0.243 0.243 0.244 0.246 0.247 0.248 0.248 0.249 0.251
+Ens2T 0.257 0.255 0.258 0.256 0.257 0.257 0.257 0.258 0.258 0.258
+Ens3T 0.249 0.247 0.250 0.248 0.250 0.250 0.251 0.251 0.251 0.252
+Ens4T 0.242 0.241 0.243 0.243 0.245 0.245 0.246 0.246 0.247 0.248
+Ens5T 0.243 0.242 0.245 0.244 0.246 0.247 0.247 0.248 0.249 0.251
+Table 6: M5 SMQL values for each method and retrain scenario combination.
 Method 7 14 21 30 60 90 120 150 180 364
@@
 Table 7: M5 CT values for each method and retrain scenario combination.
-The Tables 9 and 10 show the forecast accuracy of the different ensembles along the examined retrain scenarios for the VN1 weekly dataset, while Table 11 depicts the computing time in seconds.
 Method 1 2 3 4 6 8 10 13 26 52
-Ens2A 1.472 1.479 1.480 1.453 1.444 1.443 1.447 1.431 1.428 1.430
-Ens3A 1.517 1.520 1.504 1.462 1.451 1.435 1.450 1.429 1.429 1.430
-Ens4A 1.524 1.528 1.515 1.472 1.461 1.445 1.459 1.436 1.435 1.436
-Ens5A 1.544 1.543 1.541 1.495 1.484 1.468 1.480 1.459 1.459 1.455
-Ens2T 5.719 5.687 5.778 5.701 5.671 5.696 5.639 5.708 5.797 5.939
-Ens3T 5.536 5.492 5.556 5.400 5.326 5.369 5.112 5.246 5.497 5.782
-Ens4T 4.472 4.426 4.643 4.296 4.256 4.310 4.162 4.159 4.412 4.659
-Ens5T 3.672 3.615 3.807 3.513 3.481 3.523 3.410 3.406 3.603 3.798
+Ens2A 1.478 1.485 1.486 1.459 1.450 1.450 1.453 1.437 1.434 1.437
+Ens3A 1.523 1.526 1.510 1.468 1.458 1.441 1.456 1.435 1.435 1.437
+Ens4A 1.530 1.533 1.520 1.478 1.467 1.451 1.465 1.442 1.441 1.443
+Ens5A 1.550 1.549 1.546 1.501 1.490 1.474 1.485 1.465 1.465 1.462
+Ens2T 5.694 5.655 5.647 5.563 5.513 5.618 5.578 5.589 5.680 5.257
+Ens3T 5.286 5.297 5.262 5.263 5.290 5.253 4.943 5.118 5.454 5.503
+Ens4T 4.447 4.406 4.620 4.276 4.236 4.287 4.144 4.151 4.391 4.635
+Ens5T 3.664 3.609 3.792 3.506 3.478 3.519 3.407 3.403 3.600 3.781
 Table 8: VN1 RMSSE values for each method and retrain scenario combination.
 Method 1 2 3 4 6 8 10 13 26 52
-Ens2A 2.369 2.371 2.370 2.326 2.343 2.358 2.356 2.360 2.405 2.467
-Ens3A 2.410 2.399 2.399 2.338 2.335 2.353 2.351 2.362 2.411 2.459
-Ens4A 2.386 2.399 2.376 2.326 2.341 2.355 2.349 2.344 2.402 2.440
-Ens5A 2.375 2.371 2.383 2.318 2.334 2.341 2.357 2.346 2.405 2.446
-Ens2T 2.545 2.551 2.575 2.576 2.594 2.601 2.601 2.608 2.660 2.694
-Ens3T 2.588 2.582 2.652 2.595 2.613 2.634 2.617 2.591 2.660 2.725
-Ens4T 2.419 2.431 2.473 2.439 2.440 2.466 2.463 2.446 2.532 2.572
-Ens5T 2.339 2.335 2.355 2.341 2.361 2.380 2.391 2.380 2.456 2.492
-Table 10: VN1 MQL values for each method and retrain scenario combination.
+Ens2A 0.826 0.847 0.840 0.815 0.807 0.808 0.819 0.801 0.814 0.844
+Ens3A 0.867 0.885 0.863 0.829 0.818 0.802 0.828 0.803 0.816 0.845
+Ens4A 0.926 0.941 0.927 0.895 0.888 0.879 0.900 0.885 0.903 0.931
+Ens5A 1.013 1.024 1.014 0.990 0.982 0.975 0.989 0.971 0.999 1.018
+Ens2T 4.388 4.392 4.443 4.370 4.379 4.353 4.325 4.419 4.377 4.110
+Ens3T 3.972 4.002 4.026 4.030 4.007 3.905 3.777 3.649 4.003 4.270
+Ens4T 3.420 3.403 3.668 3.321 3.248 3.254 3.226 3.038 3.284 3.771
+Ens5T 2.832 2.805 3.011 2.738 2.687 2.691 2.673 2.513 2.725 3.105
+Table 9: VN1 SMQL values for each method and retrain scenario combination.
 Method 1 2 3 4 6 8 10 13 26 52
@@
 Table 10: VN1 CT values for each method and retrain scenario combination.
-Figures 11 and 12 show the results of the Friedman-Nemenyi test in the context of point forecast accuracy for the M5 and VN1 datasets respectively.
-Figures 13 and 14 present the test results related to the probabilistic evaluation.
+Figure 8: Friedman-Nemenyi test results based on RMSSE. The baseline scenarios are $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
+Figure 9: Friedman-Nemenyi test results based on SMQL. The baseline scenarios are $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
 Figure 10: M5 Friedman-Nemenyi test results based on RMSSE.
 Figure 11: VN1 Friedman-Nemenyi test results based on RMSSE.
-Figure 13: M5 Friedman-Nemenyi test results based on MQL.
-Figure 14: VN1 Friedman-Nemenyi test results based on MQL.
-The costs tables 12 and 13 show the estimated cost in real values of each scenario for M5 daily data.
+Figure 12: M5 Friedman-Nemenyi test results based on SMQL.
+Figure 13: VN1 Friedman-Nemenyi test results based on SMQL.
 Method 7 14 21 30 60 90 120 150 180 364 LR 390,732 208,499 150,791 112,547 69,132 52,131 48,474 46,678 38,937 33,151 XGBoost 529,679 281,358 200,778 151,635 88,200 69,495 60,082 60,155 48,189 39,861 LGBM 1,526,424 818,569 593,676 455,075 286,262 228,698 199,337 204,972 170,802 146,252 CatBoost 358,123 189,714 134,258 99,699 58,167 44,164 37,206 38,266 30,622 24,362 MLP 604,120 482,505 441,981 419,823 392,046 383,287 380,697 376,968 368,023 365,525 TCN 1,146,256 960,626 890,924 862,432 815,006 801,612 791,801 791,348 784,201 776,181 NBEATSx 729,263 561,105 514,166 466,683 425,593 411,232 400,953 401,352 393,673 387,058 NHITS 754,783 573,233 510,655 470,077 425,684 409,844 402,731 402,394 393,071 385,470 Average 754,922 509,451 429,654 379,746 320,011 300,058 290,160 290,267 278,440 269,733
@@
 Table 12: M5 estimated costs for each ensemble method and retrain scenario combination (in \$).
-The costs tables 14 and 15 show the estimated cost in real values of each scenario for VN1 weekly data.
 Method 1 2 3 4 6 8 10 13 26 52 LR 15,234 7,839 5,731 4,292 3,205 2,508 2,140 2,190 1,463 1,099 RF 1,605,768 799,597 562,896 398,954 281,246 199,214 158,959 165,975 81,529 38,865 XGBoost 34,254 17,687 12,796 9,413 6,944 5,262 4,473 4,534 2,863 2,005 LGBM 505,304 260,721 186,904 138,460 101,775 76,738 65,013 66,335 40,721 27,701 CatBoost 52,021 26,594 18,515 13,366 9,611 7,022 5,798 5,977 3,311 2,023 MLP 62,102 34,489 26,431 21,149 17,121 14,391 13,027 13,015 10,255 8,891 LSTM 82,927 49,019 38,851 31,962 27,022 23,465 21,829 21,932 18,415 16,652 TCN 72,763 41,291 31,910 25,464 20,791 17,579 15,983 16,365 12,936 11,284 NBEATSx 80,337 44,837 34,283 27,179 22,057 18,387 16,686 16,702 13,085 11,277 NHITS 80,776 45,458 34,688 27,585 22,336 18,607 16,888 16,855 13,405 11,601 Average 259,149 132,753 95,301 69,782 51,211 38,317 32,080 32,988 19,798 13,140
@@
 Table 14: VN1 estimated costs for each ensemble method and retrain scenario combination (in \$).
+Figure 14: Cost analysis.
+Comparison between base models and ensembles for the VN1 dataset.
+Each method and retrain scenario combination are shown in relative terms with respect to the baseline scenario, $r=1$.